### v1

# Causally constrained reduced-order neural models of complex turbulent dynamical systems

Abstract
We introduce a flexible framework based on response theory and score matching to suppress spurious, noncausal dependencies in reduced-order neural emulators of turbulent systems, focusing on climate dynamics as a proof-of-concept. We showcase the approach using the stochastic Charney–DeVore model as a relevant prototype for low-frequency atmospheric variability. We show that the resulting causal constraints enhance neural emulators’ ability to respond to both weak and strong external forcings, despite being trained exclusively on unforced data. The approach is broadly applicable to modeling complex turbulent dynamical systems in reduced spaces and can be readily integrated into general neural network architectures.
preprint: APS/123-QED
I Introduction
Modeling and understanding complex turbulent dynamical systems often requires abandoning full microscopic descriptions in favor of reduced-order models, where coarse-grained variables evolve according to effective stochastic dynamics [12, 68]. Such reduced stochastic descriptions are fundamental across physics, ranging from chemical physics [48, 92], neuroscience [42, 93], mathematical physics [95, 94], turbulence [67, 98], climate science [43, 65, 70, 34, 60, 51, 50], and many others. In practice, one is often faced with an inverse problem: given long trajectories of a Markovian, coarse-grained system, construct a data-driven model that faithfully reproduces its statistical and dynamical features [88, 44, 84, 74, 55, 52, 11, 102, 17, 16, 36, 3, 31, 29]. Recent advances in machine learning further contribute to inverse modeling by providing flexible neural network architectures for data-driven emulation, with promising applications in fluid dynamics [33, 76, 107], weather [82, 9], and climate science [109, 108, 40, 14, 24]. However, neural network–based models are typically agnostic to the underlying physics, which can lead to nonphysical behavior, such as violations of energy conservation - an issue that can be addressed through appropriate constraints [74, 52, 7, 78] - or incorrect responses to external forcings. The latter issue has been documented in neural weather models [105] and, especially, in climate emulators [106, 110]. This limitation highlights a central open challenge for neural emulators of physical systems: high skill in reproducing stationary statistics does not guarantee accurate responses to external perturbations, thereby limiting the use of such models for process understanding through sensitivity analysis and the exploration of “what if?” scenarios. Crucially, a model’s ability to respond to forcings is inherently linked to causality, since cause–effect relationships in physical systems are inferred through external interventions [47, 2, 61, 63], commonly realized via perturbation–response experiments (see [83, 8] in the contexts of neuroscience and climate, respectively). This connection suggests that constraining a data-driven model to respect a given “correct” causal structure should improve its ability to respond to external forcings. The resulting challenge is twofold: to infer causal relationships from stationary, unperturbed data alone, and to incorporate this information into neural emulators, all within a computationally efficient framework that is both scalable and practical. In this context, the Fluctuation–Dissipation Theorem (FDT) plays a fundamental role as it provides a bridge between forced and natural fluctuations of physical systems [53, 30, 64, 75, 39, 62]. The work of Baldovin et al. (2020) [5] showed that the FDT provides a rigorous causal inference strategy in the interventional sense directly from data, and this approach has since been adopted in a growing body of work [13, 4, 28, 27, 1, 36, 35, 29]. In this Letter, we take this line of research one step further by leveraging the FDT to constrain the training of general neural emulators of physical systems. Given a trajectory $\mathbf{x}(t)=[x_{1},x_{2},...,x_{n}](t)$ of a multivariate Markovian system, our framework proceeds in two steps: (i) use the FDT to infer the direct causal couplings $x_{j}\rightarrow x_{k}$ across degrees of freedom $x_{j}$, $x_{k}$; (ii) learn a stochastic neural model from $\mathbf{x}(t)$ while preserving the inferred causal structure. A related direction was recently proposed by Chen and Liu (2025) [16], who used causation entropy to rank candidate functions in the construction of causally constrained reduced-order models, highlighting the benefits of causally informed approaches. Our work addresses a complementary setting by providing a practical strategy to enforce a prescribed causal structure within general neural network architectures. To illustrate the methodology, we focus on climate dynamics as a representative example of a complex, turbulent dynamical system that is necessarily modeled through an effective, coarse-grained description [70, 52, 67], and for which accurate modeling of forced responses is essential [29]. We consider a stochastic version of the Charney–DeVore model [15], relevant for conceptual studies of low-frequency climate variability. This choice provides a simplified testbed that nevertheless retains key features of realistic coarse-grained climate dynamics, including strong mixing, fast decay of autocorrelations, and regular, unimodal distributions [71, 69]. We show that causally constrained neural emulators substantially improve the ability of neural networks to capture responses to both weak and strong forcings. In particular, the causal emulator accurately reproduces responses beyond the linear regime, including time-dependent shifts of the attractor in state space, despite being trained solely on stationary, unperturbed data.
II Framework: from linear response to causal constraints
Fluctuation–Dissipation Theorem (FDT) and causality
Given an nn–dimensional stochastic nonlinear system $\mathbf{x}(t)$, the FDT [64, 75] states that the linear response, in the ensemble average sense, of a degree of freedom $x_{k}(t)$ to an infinitesimally small impulse perturbation $\delta x_{j}(0)$ applied to component $x_{j}(0)$ can be written as:
$\begin{split}R_{k,j}(t)&=\frac{\delta\langle x_{k}(t)\rangle}{\delta x_{j}(0)}=-\langle x_{k}(t)s_{j}(\mathbf{x}(0))\rangle,\\ s(\mathbf{x})&=\nabla\log\rho(\mathbf{x}).\end{split}$ (1)
$\rho(\mathbf{x})$ is the system’s invariant density and $s(\mathbf{x})$ is the score function of the system [37, 35]. The focus on stochastic systems allows us to circumvent difficulties with determistic dissipative dynamical systems, where the invariant measure is singular [97, 96, 19]. The brackets $\langle\cdot\rangle$ represents ensemble averages, computed over the invariant density. Eq. (1) provides a rigorous strategy for inferring causal relations in physical systems in the interventional sense from data alone [5]. As discussed in [29], $R_{k,j}(t)$ supports multiple causal interpretations: values at the shortest time scale ($t=dt$) encode direct causal links $x_{j}\rightarrow x_{k}$. Nonzero responses at later times reflect indirect causal links. Additionally, a cumulative measure of causal influence can be defined as $D_{k,j}=\int_{0}^{\infty}R_{k,j}(t)dt$ [5]. The causal constraints proposed in this work focus on the direct causal links, $R_{k,j}(dt)$. In practice the response is computed for standardized data as $(\sigma_{j}/\sigma_{k})R_{k,j}(dt)$, $\sigma_{i}$ representing the standard deviation of $x_{i}$; for simplicity, we denote this standardized operator by $R_{k,j}$.
FDT from data
The main challenge in estimating responses via Eq. (1) lies in estimating the score function $s(\mathbf{x})$, particularly in high-dimensional settings. Traditional studies have adopted the quasi-Gaussian (qG, hereafter) ansatz [56, 64], in which $\rho$ is approximated by a Gaussian. The qG approximation has proven effective for predicting mean responses, even for nonlinear problems [64, 39, 71, 35]. Recent developments in score-based generative modeling [100, 99], offer new ways forward to overcome previous approximations by parametrizing the score function directly from data. Importantly, when provided with enough data, these new strategies scale to high-dimensional systems such as quasi-geostrophic flows [85], idelized ocean flows [101], weather [91], climate [10], fluid dynamics [57], among others. The estimated score $s(\mathbf{x})$ can then be inserted into the analytical response formula in Eq. (1), circumventing the need for explicit approximations of the invariant density, as in the quasi-Gaussian case. This approach was first proposed in [37]. These advances are leading to a set of score-estimation techniques rather than a single strategy. For example, [35] highlighted different denoising score-matching procedures tailored to the dimensionality of the underlying system. Here, we further contribute to this “FDT–generative modeling toolbox”. Specifically, we adopt the score-matching estimator proposed in [46], allowing us to infer the score directly on the attractor, rather than up to a small noise perturbation (e.g., as in denoising score matching). The estimated score is used as an initial approximation and is then refined using the correction procedure introduced in [37]. Details are reported in Section I of the Supplemental Material.
Causal constraints.
Consider a stochastic dynamical system
$\dot{x_{i}}(t)=f_{i}(x_{1},x_{2},...,x_{n})+\sigma_{i}\xi_{i}(t),~\text{with}~i=1,2,...,n,$
where $\sigma_{i}\xi_{i}(t)$ is a stochastic forcing, $\bm{\xi}(t)$ being a standard Gaussian white noise process. A perturbation to $x_{j}$ can affect $x_{k}$ at the next time step only if $f_{k}$ depends on $x_{j}$. Thus, in the limit $dt\to 0$, $\partial f_{k}/\partial x_{j}\equiv 0$ implies the absence of a direct causal link and consequently, a vanishing short-time response $R_{k,j}(dt)=0$. In practice, the system’s Jacobian is not observable from data, and a reliable numerical estimation is unfeasible in high-dimensional systems. However, the short-time response $R_{k,j}(dt)$ can be efficiently estimated from unperturbed data using the FDT in Eq. (1) as shown in [35]. A vanishing response at the shortest resolved time scale thus implies $\big\langle\partial f_{k}(\mathbf{x})/\partial x_{j}\big\rangle=0$ for $j\neq k$. This information on vanishing entries of the ensemble-averaged Jacobian is used to constrain neural architectures. Thus, given a long trajectory of $\mathbf{x}(t)$, we first estimate the response operator $\mathbf{R}(dt)\in\mathbb{R}^{n,n}$ with the score-matching procedure introduced earlier. We then proceed by constructing a binary adjacency matrix $\mathbf{A}\in\mathbb{R}^{n,n}$, by thresholding $\mathbf{R}(dt)$. $A_{k,j}=0$ denoting the absence of a direct causal link $x_{j}\rightarrow x_{k}$, and $A_{k,j}=1$ denoting its presence. One way to threshold $R_{k,j}(dt)$ is to examine its values on a logarithmic scale; a visual inspection of the response magnitudes often reveals a clear separation between significant links and noise. The absence of such separation may itself be informative, suggesting a densely connected system, in which case no additional causal constraints are needed. To provide an unsupervised criterion, we propose the following procedure:
  1. i.
    Self-interactions. We assume that for any dissipative dynamical system $x_{j}\rightarrow x_{j}$ and therefore $A_{j,j}=1$.
  2. ii.
    Cross-interactions. We consider the off-diagonal values of $R_{k,j}(dt),j\neq k$, and apply a logarithmic transformation $\ln|R_{k,j}(dt)|$. This helps separate potential causal links (larger responses) from the bulk of spurious, near-zero interactions. A $k$-means algorithm [86] ($k=2$) is then used to partition these log-scaled values into two clusters, corresponding to “significant” and “insignificant” responses. An entry of $A_{k,j}=1$ is set if its corresponding $\ln|R_{k,j}(dt)|$ value belongs to the cluster with the larger centroid; otherwise, it is set to 0.
The causal Adjacency matrix is then used to constrain the training of a neural emulator. Given a $n$-dimensional trajectory $\mathbf{x}(t)=[x_{1},x_{2},...,x_{n}](t)$ we fit a neural operator $\mathbf{f}$ such that $\dot{\mathbf{x}}(t)=\mathbf{f}(\mathbf{x})+\mathbf{\Sigma}\bm{\xi}(t)$. The causal constraints are imposed on the deterministic drift $\mathbf{f}$. We learn $\mathbf{f}$ via a mean squared error (MSE) minimization such that $\Delta\mathbf{x}=\mathbf{x}(t+1)-\mathbf{x}(t)=\mathbf{f}(\mathbf{x}(t))$, where the time unit has been rescaled by the time sampling interval. Details on the learning process are found in the SM, Section III. A “vanilla”, unconstrained operator $\mathbf{f}$ is trained by minimizing MSE as $\mathcal{L}=MSE(\Delta\mathbf{x},\mathbf{f}(\mathbf{x}))$. The causally constrained emulator is trained by minimizing the following loss function:
$\mathcal{L}=MSE(\Delta\mathbf{x},\mathbf{f}(\mathbf{x}(t)))+\lambda\sum_{k,j\in\mathcal{S}}(\frac{\partial f_{k}}{\partial x_{j}})^{2}~,$ (2)
where $\mathcal{S}$ is the set of pairs $(k,j)$ for which the causal adjacency matrix $A_{k,j}=0$. $\lambda$ is a parameter determining the strength of the regularization. The derivatives are obtained via automatic differentiation [81]. We adopt soft constraints for two main reasons. First, they are computationally efficient [78]. Second, in real-world, high-dimensional climate systems with partial observations and finite data, causal estimates are inherently uncertain. In this regime, hard constraints based on imperfect causal information will introduce biases into the learned dynamics. Soft constraints, in contrast, provide a flexible form of regularization, ensuring an appropriate balance between data fidelity, as captured by the MSE, and causal regularization. Additionally, we note that the regularization in Eq. (10) can be adapted to novel architectures, including generative models [18], which we leave to future work.
Possible limitations for idealized systems.
The proposed constraints leverage information about zeros of the ensemble-averaged Jacobian to constrain the state-dependent Jacobian. While this choice is practical and efficient, it can be limiting in idealized systems, for example because of exact symmetries. In such cases, nonzero local couplings may cancel in ensemble averages, yielding $\langle\partial f_{k}(\mathbf{x})/\partial x_{j}\rangle=0$ even if $\partial f_{k}(\mathbf{x})/\partial x_{j}\neq 0$, as occurs, for example, in the Lorenz ’63 model [58]. In such idealized settings, applying the proposed approach may require a symmetry-reduction step, as advocated in [22] and demonstrated in [77] for the Lorenz ’63 system. These limitations are not restrictive for realistic turbulent system considered here, such as weather and climate, which lack exact symmetries, are only partially observed, and admit multiple effective coarse-grained representations. In this regime, dominant interactions among coarse-grained variables are often well approximated by a state-independent linear operator, with nonlinearities providing higher-order corrections [88, 87, 89, 90, 111, 41, 64, 70]. There is no unique law to be discovered, instead the learned drift $\mathbf{f}$ represents one admissible effective realization, regularized to ensure physically consistent response behavior and agreement with observed statistics. Further details are provided in Section II of the SM.
Practical assessment of causal constraints from data alone.
Finally, the utility of causal constraints for a given system can be readily assessed by fitting a constrained model and an unconstrained counterpart. A key diagnostic is the comparison of the response operator computed from the emulators (see Section V of the SM) with the FDT-based operator in Eq. (1). When the inferred causal structure is informative, the constrained emulator improves the representation of linear responses without degrading the representation of stationary statistics. Moreover, comparing modeled and FDT-inferred response operators provides a natural criterion for a posteriori calibration of neural emulators based on forced responses. Related response-based calibration ideas have been explored in previous studies [72, 73]; a systematic formulation tailored to neural emulators and causal constraints is left for future work.
III Causal constraints in practice
Test-model for low-frequency climate variability.
To showcase the framework, we consider a stochastic version of the Charney–DeVore model [15], “CdV” hereafter. The model was obtained as a six-dimensional Galerkin truncation of the barotropic vorticity equation on a $\beta$-plane channel with topography. The formulation considered here is the one further proposed by De Swart [23]. The model has been widely studied in the context of large-scale atmospheric circulation [20, 21, 26], as well as in computational studies of regime transitions and rare events [38]. The model takes the following form:
$\dot{x_{1}}$ $=\tilde{\gamma}_{1}x_{3}-C(x_{1}-x_{1}^{*})+\sigma\xi_{1}(t),$ (3)
$\dot{x_{2}}$ $=-(\alpha_{1}x_{1}-\beta_{1})x_{3}-Cx_{2}-\delta_{1}x_{4}x_{6}+\sigma\xi_{2}(t),$
$\dot{x_{3}}$ $=(\alpha_{1}x_{1}-\beta_{1})x_{2}-\gamma_{1}x_{1}-Cx_{3}+\delta_{1}x_{4}x_{5}+\sigma\xi_{3}(t),$
$\dot{x_{4}}$ $=\tilde{\gamma}_{2}x_{6}-C(x_{4}-x_{4}^{*})+\varepsilon(x_{2}x_{6}-x_{3}x_{5})+\sigma\xi_{4}(t),$
$\dot{x_{5}}$ $=-(\alpha_{2}x_{1}-\beta_{2})x_{6}-Cx_{5}-\delta_{2}x_{4}x_{3}+\sigma\xi_{5}(t),$
$\dot{x_{6}}$ $=(\alpha_{2}x_{1}-\beta_{2})x_{5}-\gamma_{2}x_{4}-Cx_{6}+\delta_{2}x_{4}x_{2}+\sigma\xi_{6}(t).$
In what follows, the six variables $x_{i}$ serve as an effective reduced-order description of an underlying high-dimensional dynamical system, from which we aim to learn a causally constrained neural emulator. The parameters and their physical interpretation are reported in Section IV of the SM. $\xi_{i}(t),i=1,...,6$ is a Gaussian white noise with zero mean, unit variance, and $\sigma$ is the noise amplitude. This stochastic forcing mimics the effect of unresolved fast processes. The model is an ideal first test case for our framework of large-scale, low-frequency climate dynamics for two main reasons. First, Eq. (3) features the abstract form $\dot{\mathbf{x}}=\mathbf{F}+\mathbf{L}\mathbf{x}+\mathbf{B}(\mathbf{x},\mathbf{x})+\mathbf{\Sigma}\bm{\xi}(t)$, with energy conserving quadratic nonlinearities such that $\mathbf{x}\cdot\mathbf{B}(\mathbf{x},\mathbf{x})=0$; a structure that naturally arises when geophysical turbulent flows are projected onto orthogonal bases [64, 65, 66, 67]. Second, this system is strongly mixing, with fast decay of autocorrelations and regular, unimodal probability distributions: such properties are relevant for coarse-grained climate dynamics as found both in models and observational data [32, 6, 104, 69], and underlie the reduced-order modeling of realistic, complex turbulent systems [68]. In this respect, the resulting dynamics differ qualitatively from those of prototypical chaotic models like the Lorenz ’63 system [58], which is weakly mixing, possesses exact symmetries, and exhibits a singular invariant measure [69].
We simulate this model for $T=5\cdot 10^{7}$ time-steps at $dt=0.01$, i.e. $5\cdot 10^{5}$ Model Time Units (MTU). This long integration serves as a stationary run to (i) infer the response operators $R_{k,j}(t)$ from data alone and (ii) to build neural emulators. Two emulators, with and without causal constraints, are then compared in terms of their ability to reproduce both stationary and perturbed statistics. While a comprehensive analysis is provided in the Supplemental Material, the main text highlights the identification of causal links and the ability of the causally constrained emulator to capture time-dependent responses in ensemble mean and variance under strong forcings.
Causal Adjacency matrix from data.
Given a long trajectory $\mathbf{x}(t)=[x_{1},...,x_{6}](t)$ of the CdV model, we infer the direct causal links through the steps proposed in the previous Section. The strategy is highlighted step by step in Figure 1. We infer the response operator at the shortest time scale $R_{k,j}(dt)=-\langle x_{k}(dt)s_{j}(\mathbf{x}(0))\rangle$ via score-matching, consider only the off-diagonal elements, and apply the transformation $\ln|R_{k,j}(dt)|$; see Figure 1(a). We then apply a $k$-means algorithm ($k=2$) to split the large responses from the spurious non-causal ones; see Figure 1(b). This information is used to define the Adjacency matrix in Figure 1(c), thus defining the causal functional form of the system to be imposed in the neural emulator; see 1(d). In Section VII of the SM we also report results for the inference of the response operator $R_{k,j}(t)$ over multiple time lags $t$ using both the score-matching and quasi-Gaussian approximations, further illustrating the advantages of the score-matching approach.
Figure 1: Causal constraining of the CdV model (Eq. (3)) from data through the FDT. Panel (a): The 30 off-diagonal responses $\ln|R_{k,j}(dt)|$ (evaluated at the shortest time scale $t=dt$), flattened and plotted against a dummy index. Panel (b): Unsupervised identification of a response threshold via $k$-means clustering ($k=2$): clusters are colored in black and blue. Panel (c): Causal adjacency matrix: $A_{k,j}=1$ is assigned if its corresponding response value lies in the higher centroid cluster; otherwise $A_{k,j}=0$. Panel (d): “causal” deterministic functional form to be enforced in the emulator loss function in Eq. (10).
Neural emulators.
We learn two stochastic neural models from the long trajectory $\mathbf{x}(t)=[x_{1},...,x_{6}](t)$ of the CdV model: one with causal constraints and one without. The latter referred to as “vanilla” emulator. The causally constrained emulator leverages the Adjacency matrix $A_{k,j}$ inferred in Figure 1(c) to suppress spurious, non-causal terms through the new loss function in Eq. (10). The set $\mathcal{S}$ of pairs $(k,j)$ to be used in Eq. (10) is given by the zeros of $A_{k,j}$ in Figure 1(c) and equal to $\mathcal{S}=\{(1,2),(1,4),(1,5),(1,6),(2,5),(3,6),(4,1),(5,2),(6,3)\}$. Both neural emulators have been trained with the same hyper-parameters, the only difference being the addition of the regularization term as in Eq. (10) for the causally constrained model.
Stationary statistics.
Given the two emulators defined above, we simulate a long trajectory of $T=10^{7}$ time steps. We then examine two stationary statistics: the invariant probability distribution (PDF) and the autocorrelation functions (ACFs). The analysis is reported in Section VI of the SM: both the vanilla and the causally constrained emulator reproduce these stationary statistics well.
Perturbed statistics: linear regime.
We assess the ability of the emulators to reproduce perturbed statistics in the linear-response regime by analyzing the impulse response operator for the ensemble mean and variance, following [29]. A small impulse perturbation $\delta x_{j}(0)$ is applied at $t=0$, and the response of observables $A(x_{k}(t))=x_{k}(t)$ (mean) and $A(x_{k}(t))=(x_{k}(t)-\mu_{k}(t))^{2}$ (variance) is measured, $\bm{\mu}(t)$ representing the time-dependent mean. Here, we perturb $x_{1}(0)$ and report the self-response of $x_{1}(t)$ in mean and variance (Fig. 2). The vanilla emulator reproduces the response operator accurately, as expected given the large training dataset and Markovianity of the system [29]. However, the causally constrained emulator yields a systematically improved representation of both mean and variance responses. This improvement persists for the full response operator $R_{k,j}(t)$ across all variables (see Figures 4–5 of the Supplemental Material). A quantitative summary of this improvement is provided by the total time-dependent MSE, aggregated over all $j$ and $k$, between the emulators and the ground truth, reported in Figure 6 of the SM.
Figure 2: Perturbed statistics: linear regime. Self-response operator of mean and variance for the variable $x_{1}$. Panel (a): Time-dependent, response in ensemble mean of variable $x_{1}(t)$ to an impulse perturbation $\delta x_{1}(0)$ imposed on $x_{1}(0)$ at time $t=0$. Formally: $R_{1,1}(t)=\delta\langle x_{1}(t)\rangle/\delta x_{1}(0)$. Panel (b) same as Panel (a) but for response in ensemble variance. Responses are computed using an ensemble of $N_{e}=10^{6}$ members. The magnitude of the impulse perturbation is set to $10^{-1}\sigma_{1}$, $\sigma_{1}$ being the standard deviation of $x_{1}$. Labels: “Numerical” (numerical model), Model vanilla” (vanilla, unconstrained, emulator), and “Model causal” (causally constrained emulator).
Perturbed statistics: nonlinear regime.
We now probe the nonlinear response regime by applying a large step forcing $\mathbf{F}$ to both the numerical model and the neural emulators. As in the linear response analysis, the forcing is applied to $x_{1}$. The forced dynamics are given by $\dot{\mathbf{x}}=\mathbf{F}+\mathbf{f}(\mathbf{x})+\mathbf{\Sigma}\bm{\xi}(t)$, where $\mathbf{F}=(\sigma_{1},0,0,0,0,0)$ for $t\geq 0$, $\sigma_{1}$ being the std. dev. of the $x_{1}$ variable. The forcing amplitude is intentionally set very high to induce a strongly nonlinear response. We simulate an ensemble of $10^{5}$ forced realizations and measure the time-dependent response of the ensemble mean and variance relative to the unperturbed stationary distribution. The forcing alters the system’s attractor through time-dependent feedbacks across degrees of freedom. Figure 3 shows the resulting self-response of the marginal distribution of $x_{1}(t)$; responses of the full system are reported in Figures 7–8 of the Supplemental Material. The numerical model exhibits a pronounced transient response, relaxing toward a new equilibrium after approximately $40$ model time units. Therefore, the external forcing modifies the characteristics of the attractor in state space. The causally constrained emulator accurately reproduces these nonlinear forced responses, particularly in the ensemble mean. By contrast, the vanilla emulator becomes unstable after a short transient under the same forcing. This behavior is robust: when large step forcings are applied to each degree of freedom, the vanilla emulator becomes unstable in three out of six cases, whereas the causally constrained emulator remains stable for all forcings (see Section VIII of the SM). In general, despite the large forcing amplitude, we observe that the causally constrained emulator provides a good approximation of the forced responses in both ensemble mean and variance.
Figure 3: Perturbed statistics: nonlinear regime. Time-dependent response of the ensemble mean and variance of $x_{1}$ to a strong step function forcing. Forcing: step function $\mathbf{F}=(\sigma_{1},0,0,0,0,0)$ for $t\geq 0$, applied on the right-hand side of the numerical model and emulators. $\sigma_{1}$ is the standard deviation of $x_{1}$. Responses are computed using an ensemble of $N_{e}=10^{5}$ members. Panel (a): response in ensemble mean. Panel (b): response in ensemble variance. Labels: “Numerical” (numerical model), “Model vanilla” (vanilla, unconstrained, emulator), and “Model causal” (causally constrained emulator).
IV Discussion and conclusions
We introduced a strategy that leverages the Fluctuation–Dissipation Theorem and score matching to (i) identify direct causal links from long trajectories of Markovian physical systems and (ii) impose this causal information as constraints in neural emulators. These constraints improve the ability of reduced-order neural emulators to capture responses to both weak and strong external perturbations, even if trained on unforced data alone. In practice, the choice of a Markovian, coarse-grained representation of an underlying high-dimensional dynamical system is itself a central challenge, as emphasized in classical work [80] and revisited in recent studies [73, 63, 29]. However, given an appropriate choice of coarse-grained variables, the proposed constraints can be readily implemented in general neural network architectures. The framework is modular and can be combined with any data-driven estimator of the FDT operator, including recent proposals such as [62, 25], as well as with alternative causal discovery methods, for example [79]. A natural next step is to extend this framework to observational climate data and to develop causally consistent, reduced-order neural models of the global atmosphere and ocean, including within emerging probabilistic architectures such as stochastic interpolants [18, 112, 78]. This perspective also opens the possibility of combining causal constraints with complementary physical constraints, such as energy conservation [7], to jointly enforce stability and physical fidelity in neural emulators. Additionally, in realistic applications, data are often scarce and partially observed, leading to imperfect causal estimates. An important open direction is to assess how accurate such estimates must be - and whether approximate, quasi-Gaussian FDT operators may already suffice [69] - to construct imperfect emulators that nevertheless exhibit skill in responding to external perturbations. Finally, while motivated by climate dynamics, the proposed strategy applies to neural modeling of general complex turbulent dynamical systems in reduced coordinate spaces. In such settings, causal constraints provide a practical and readily implementable addition to general neural network architectures.
Code availability
Acknowledgements.
This research was supported by Schmidt Sciences, LLC, through the M2LInES project. Computational resources and support were provided in part by the NYU IT High Performance Computing facilities, services, and staff. F.F. is grateful to Rory Basinski and Matthieu Blanke for many enriching discussions on this subject.
Supplemental Material for:
Causally constrained reduced-order neural climate emulators via response theory
Fabrizio Falasca and Laure Zanna
Courant Institute School of Mathematics, Computing and Data Science, New York University
(Dated: )
Fabrizio Falasca Laure Zanna
Figure 1: Schematic of the proposed framework. Step 1: From a long trajectory of a Markovian system, the Fluctuation–Dissipation Theorem is used to infer the system’s causal structure. Step 2: This inferred structure is then imposed as a soft constraint during neural-network training, yielding causally constrained emulators. The example shows time series from the Charney–DeVore model (Eq. (3) in the main text) and the discovered causal graph, but the approach applies to general stochastic dynamical systems.
I Fluctuation-Dissipation Theorem (FDT) from data: quasi-Gaussian and score matching
The FDT.
Consider a $n$-dimensional system $\mathbf{x}(t)=[x_{1},x_{2},...,x_{n}](t)$. The Fluctuation-Dissipation Theorem (FDT) states that the time-dependent response of an observable $A(x_{k}(t))$ to a small impulse perturbation $\delta x_{j}(0)$ imposed on $x_{j}(0)$ at time $t=0$, can be retrieved solely from stationary statistics of the system as:
$R_{k,j}(t)=\lim_{\delta x_{j}(0)\to 0}\frac{\delta\langle A(x_{k}(t))\rangle}{\delta x_{j}(0)}=-\Big\langle x_{k}(t)\frac{\partial\ln\rho(\mathbf{x})}{\partial x_{j}}\Big|_{\mathbf{x}(0)}\Big\rangle.$ (1)
where $\rho(\mathbf{x})$ represents the invariant probability distribution of the system [30, 75]. The FDT establishes a link between perturbed quantities and stationary statistics, i.e. the invariant measure of the system [64, 75]. The brackets $\langle\cdot\rangle$ represents ensemble averages, in practice computed using temporal averages under the assumption of ergodicity. Application of the FDT in the main text focused on the observable $A(x_{k}(t))=x_{k}(t)$. In this case, Eq. 1 represents the impulse response in the ensemble mean.
I.1 The FDT from data
Quasi-Gaussian (qG) approximation.
A long standing problem in inferring the FDT response operator is the need to estimate the gradients of logarithm of the invariant distribution $s(\mathbf{x})=\nabla\ln\rho(\mathbf{x})$, the so-called score function [46]. Traditionally, the quasi-Gaussian (qG) approximation has been used [56, 64], by replacing $\rho(\mathbf{x})$ by a Gaussian distribution $\rho^{G}(\mathbf{x})$ with the same mean and second moments. In this case the score can be directly is:
$s^{G}(\mathbf{x})=\nabla\ln\rho^{G}(\mathbf{x})=-\mathbf{\Sigma}^{-1}(\mathbf{x}-\bm{\mu}).$ (2)
$\bm{\mu}\in\mathbb{R}^{n}$ and $\mathbf{\Sigma}\in\mathbb{R}^{n,n}$ are the mean and covariance matrix of the data, respectively. Note, that while a Gaussian approximation is used for $\rho(\mathbf{x})$, the response formula in Eq. (1) is still computed over the original invariant measure [64, 35]. In other words, this approximation differs from linear inverse model strategies which assume a Gaussian measure from the outset. The qG approximation has high skill in predicting responses in the ensemble mean even in nonlinear systems [39, 64, 5] but it can introduce biases for higher order moments [71, 35].
Score-based generative modeling and FDT.
Recent advances in generative modeling provide novel tools for estimating the score function directly from data, overcoming assumptions such as the qG approximation. In particular, [100] showed that score estimation can be recasted as a denoising problem. This led to scalable diffusion models that integrate a reverse-time Langevin equation from a reference distribution (typically Gaussian) toward the data distribution [85]. As in the FDT framework, a central object of these frameworks is the score function. Giorgini et al. (2024) [37], first proposed to use denoising score matching strategies in the FDT context. In practice, the true score on the system’s attractor can be then approximated by the score of a distribution perturbed with a small amount of noise [37]. This strategy has been successfully applied to stochastic dynamical systems of increasing complexity [37, 35]. The broader appeal of this approach is its connection between established ideas from statistical physics (FDT) and modern generative modeling.
Score matching.
In this study, we extended the prior work in [37, 35] by learning the score directly on the attractor using the classic result of Hyvärinen (2005) [46], hereafter “score matching”. Given $T$ samples of a $n$ dimensional dynamical system $\mathbf{x}\in\mathbb{R}^{T,n}$, we parameterize the score function $s_{\theta}(\mathbf{x})$ with a neural network $s$ and parameters $\theta$. Specifically, $s_{\theta}$ is approximated with a multilayer perceptron (MLP). The network is trained by minimizing the following loss:
[…]
where the expectation $\mathbb{E}_{\rho(x)}$ is approximated by a time average over the available samples. The specific implementation used in this work relies on an approximation similar to the sliced score-matching approach proposed in [99]. Specifically, we considered the Hutchinson Trace Estimation [45] to compute the trace in Eq. (3). As in Giorgini et al. (2025) [35], in the case of high-dimensional dynamics we suggest using the common denoising score matching implemented with a U-Net architecture.
Practical considerations and correction procedure.
The score-matching procedure introduced by Hyvärinen (2005) [46] provides an elegant way to estimate the score $s(\mathbf{x})$ directly on the system’s inertial manifold. In practice, however, this estimation can be prone to several sources of error. First, ensemble averages must be replaced by empirical averages computed from finite datasets. Second, the inferred score inevitably depends on the chosen neural-network architecture and its associated hyperparameters. Nevertheless, the theoretical formalism of the FDT, provides analytical constraints on the instantaneous response $\mathbf{R}(t=0)$. As shown in [37], these constraints can be used as a posteriori correction of the estimated score, thereby ensuring reliable results. Such corrections are non-unique, but they typically produce only small adjustments to the inferred response and thus eliminate the need for more sophisticated architectures. In what follows, we detail the analytical correction for score-based estimates relevant to the ensemble-mean response. We denote the ideal score by $s(\mathbf{x})$, the score inferred through score matching by $s_{\epsilon}(\mathbf{x})$, and the corrected score by $s_{c}(\mathbf{x})$. The system’s dynamics is encoded in a data matrix $\mathbf{x}\in\mathbb{R}^{T,n}$, where $T$ is the number of temporal samples and $n$ the system’s dimensionality. The correction procedure is dependent on the type of response analyzed. If the interest is in the response in ensemble mean, i.e. observable $A(x_{k}(t))=x_{k}(t)$ in Eq. (1), then it is possible to use the correction already outlined in [37]. The response operator in the ensemble mean at time $t=0$ of any system is:
$\mathbf{R}(0)=\mathbf{I}$ (4)
$\mathbf{I}$ being the identity matrix. The empirical estimation of the instantaneous response operator with the inferred score $s_{\epsilon}(\mathbf{x})$ will lead to small errors, so that
$\mathbf{R}_{\epsilon}(0)=-\frac{1}{T}\mathbf{x}^{\text{T}}\cdot s_{\epsilon}(\mathbf{x})=\mathbf{I}-\boldsymbol{\epsilon}$ (5)
where $\mathbf{R}_{\epsilon}(0)$ denotes the estimated response. It is then possible to correct the inferred score as:
$s_{c}(\mathbf{x})=s_{\epsilon}(\mathbf{x})\cdot(\mathbf{I}-\boldsymbol{\epsilon})^{-1}.$ (6)
The corrected score $s_{c}(\mathbf{x})$ defined in Eq. (6) guarantees the equivalence Eq. (4) and it is the one used in the computation of responses in ensemble mean in the main text. Note that while this correction is non-unique, different versions lead to small differences as argued in the SM of [36].
Response operator via score matching: response in the ensemble mean.
Given the proposed correction in Eq. (6), we now infer the response operator for the Charney-DeVore system in Eq. (3) in the main text. We compare the prediction of the qG approximation and the score-matching estimation against the the numerical ground truth and show the results in Figure 2. First, we note that the quasi-Gaussian approximation leads to good, first-order results even in a strongly nonlinear model such as the Charney–DeVore model. The FDT inferred through score matching is generally more accurate than the quasi-Gaussian approximation.
Figure 2: Time-dependent, responses in ensemble mean $x_{k}(t)$ to an impulse perturbation in the CdV model (Eq. (3) in the main text) as predicted by the FDT using the quasi-Gaussian and score-matching approximations. Example: column (1), row (2) quantifies the time-dependent mean response of $x_{k=2}(t)$ given a small, impulse perturbation imposed on $x_{j=1}(0)$ at time $t=0$.
II Possible limitations of the causal constraints and practical considerations
Causal relations in high-dimensional stochastic dynamical systems are most naturally, and practically, identified at the level of ensemble-averaged quantities [29]. Specifically, the response operator at the shortest time-scale can be written in terms of the ensemble averaged Jacobian as $\mathbf{R}(dt)=\langle\mathbf{J}(\mathbf{x})\rangle dt+\mathbf{I}$, where $\mathbf{J}(\mathbf{x})$ denotes the Jacobian, $\mathbf{I}$ is the identity matrix, and $dt\rightarrow 0$. Accordingly, the causal constraints introduced in the main text, exploit vanishing entries of the ensemble-averaged Jacobian, $\langle\partial f_{k}(\mathbf{x})/\partial x_{j}\rangle=0$, to suppress the corresponding state-dependent couplings $\partial f_{k}(\mathbf{x})/\partial x_{j}$ in the neural network loss. In other words, the central assumption of this work is that knowledge of the zeros of the ensemble averaged Jacobian already provides an informative constraint for guiding the architecture of general neural emulators. A complication arises in idealized systems with exact symmetries, such as the Lorenz ’63 and Lorenz ’96 systems [58, 59]. In such cases $R_{k,j}(dt)=0$ can arise from exact cancellations in the ensemble average, even though $\partial f_{k}(\mathbf{x})/\partial x_{j}\neq 0$ pointwise. In such cases, a vanishing short-time response does not necessarily imply the absence of an instantaneous coupling. One possible strategy in such idealized settings is to explicitly reduce the symmetries of the underlying dynamical system, following symmetry-reduction approaches as advocated in [22]. For example, in the Lorenz ’63 system [58], this corresponds to considering the so-called “proto-Lorenz” system [77], in which the symmetries of the original system are reduced. More importantly, however, we argue that this limitation does not constitute a practical obstacle for reduced-order modeling of real-world turbulent climate systems, which are the primary focus of this work. Such systems typically lack exact symmetries and admit no unique coarse-grained representation. The effective drift term $\mathbf{f}$ is validated on its ability to reproduce stationary and statistics given a specific choice of coarse-grained representation. In this context, we find it useful to provide additional qualitative justification for why focusing on zeros of the ensemble-averaged Jacobian is not a fundamental limitation. First, we discuss a qualitative mathematical argument in the context of linear inverse models of climate. Second, we emphasize a more general point regarding the non-uniqueness of effective dynamics in data-driven modeling of coarse-grained systems.
  • In the context of reduced-order modeling of large-scale climate variability, typically constructed from anomalies, there is extensive evidence that linear inverse models (LIMs), i.e. $d\mathbf{x}=\mathbf{L}\mathbf{x}dt+\mathbf{\Sigma}d\mathbf{W}$, as originally proposed by Penland [88], provide a useful effective approximation of the dynamics and serve as powerful forecasting tools; see [88, 87, 89, 90, 111, 41] among many others. A natural generalization incorporates quadratic nonlinearities,
    $d\mathbf{x}=(\mathbf{L}\mathbf{x}+\mathbf{B}(\mathbf{x},\mathbf{x}))dt+\mathbf{\Sigma}d\mathbf{W}$
    as discussed, for example, in [64]. In what follows, we focus on nonlinear extensions of the LIM within this quadratic framework. In this case, the $k$-th component of the drift can be written as
    $f_{k}(\mathbf{x})=\sum_{m}L_{k,m}x_{m}+\sum_{m,n}B_{k,m,n}x_{m}x_{n}~.$
    and the corresponding Jacobian entry takes the form
    $J_{k,j}(\mathbf{x})=\partial f_{k}(\mathbf{x})/\partial x_{j}=L_{k,j}+\sum_{m}B_{k,j,m}x_{m}+\sum_{n}B_{k,n,j}x_{n}~.$
    Taking the ensemble average yields
    $\langle J_{k,j}(\mathbf{x})\rangle=\langle\partial f_{k}(\mathbf{x})/\partial x_{j}\rangle=L_{k,j}+\sum_{m}B_{k,j,m}\langle x_{m}\rangle+\sum_{n}B_{k,n,j}\langle x_{n}\rangle~.$
    In the common modeling regime where climate variability is represented in terms of anomalies, one typically has $\langle x_{n}\rangle=0$ for all components, leading to
    $\langle J_{k,j}(\mathbf{x})\rangle=L_{k,j}~.$
    In this modeling setting, the ensemble-averaged Jacobian therefore coincides with the linear operator $\mathbf{L}$. The empirical success of linear inverse models for coarse-grained climate variability suggests that, at this level of coarse-grained description (typically around 20 principal components of a given climate field, see e.g. [87]), the linear operator $\mathbf{L}$ approximates the dominant direct couplings, with nonlinear terms acting as corrections. This linear operator is obviously state independent. As a result, in this example, any vanishing entry $\langle J_{k,j}\rangle=L_{k,j}=0$ reflects the absence of a linear coupling, rather than a cancellation induced by ensemble averaging, therefore providing intuition for why vanishing entries of the ensemble-averaged Jacobian are structurally meaningful constraints.
  • More broadly, this work highlights a principle relevant to data-driven modeling of realistic multiscale systems. As discussed in [29], the effective drift $\mathbf{f}$ inferred from data is inherently non-unique and depends on modeling choices such as variable selection, resolution, filtering, and projection onto latent representations. In this setting, there is no single “true” dynamics; instead, models are evaluated by their ability to reproduce stationary and perturbed statistics at the chosen level of coarse graining. From this perspective, a vanishing short-time response $R_{k,j}(dt)=0$ provides sufficient justification to construct an effective model
    $d\mathbf{x}=\mathbf{f}(\mathbf{x})dt+\mathbf{\Sigma}(\mathbf{x})d\mathbf{W}$
    in which the structural constraint $\partial f_{k}(\mathbf{x})/\partial x_{j}=0$ is explicitly enforced. The validity of such constraints is then assessed by the accuracy of the resulting stationary and forced responses, rather than by comparison to a non-existent unique underlying dynamics.
Assessment of causal constraints.
Finally, causally constrained emulators can be readily compared with an unconstrained counterpart by evaluating their representation of the linear response operator against the FDT-based estimate in Eq. (1) of the main text. Such comparisons should be performed at short time lags in order to avoid spurious responses that may arise at longer lags when using the FDT [28]. The causal constraints introduced here should be applied when they improve the representation of the response operator relative to the unconstrained model. In systems with exact symmetries, as discussed above, any biases introduced by the constraints will already be apparent in the representation of stationary variability.
III Details on the emulator formulations
We report the main technical details for fitting the emulators in the main text. This Section follows Appendix C in [29] and we refer to that work for more details. Consider a $n$-dimensional dynamical system described by a stochastic differential equation $d\mathbf{x}=\mathbf{f}(\mathbf{x})dt+\mathbf{\Sigma}d\mathbf{W}$. This equation can be discretized by the forward Euler method as follows:
[…]
where $\bm{\xi}(t)$ is a Gaussian white noise with zero mean, unit variance, and $\sqrt{\Delta t}\bm{\xi}(t)$ is the increment of the Wiener process $\Delta\mathbf{W}=\mathbf{W}(t+\Delta t)-\mathbf{W}(t)$. Without loss of generality, we rescale time units by the time sampling interval such that a non-dimensional time increment is $\Delta t=1$ [55, 54, 52, 103]. The strategy is then to fit a rescaled version of Eq. (7) from observational data alone. Such rescaled equation takes the following form:
$\begin{split}\Delta\mathbf{x}=\mathbf{x}(t+1)-\mathbf{x}(t)=\mathbf{f}(\mathbf{x}(t))+\mathbf{\Sigma}\bm{\xi}(t).\end{split}$ (8)
In what follows, we will not distinguish between the original model terms (e.g., $\mathbf{f}$) and their emulator counterparts, in order to avoid unnecessary notation and maintain a clear and concise presentation. We explicitly clarify the distinction whenever it becomes relevant. We proceed in two steps: fitting the determistic drift and then the stochastic noise. We first focus on fitting the drift. We compute $\Delta\mathbf{x}=\mathbf{x}(t+1)-\mathbf{x}(t)$ and fit the deterministic dynamics such that $\Delta\mathbf{x}\approx\mathbf{f}(\mathbf{x}(t))$. In the main paper, we considered two cases:
  1. i.
    Unconstrained drift. The vanilla emulator is trained to minimize the MSE between $\Delta\mathbf{x}$ and $\mathbf{f}(\mathbf{x}(t))$ as:
    $\mathcal{L}=MSE(\Delta\mathbf{x},\mathbf{f}(\mathbf{x}))~.$ (9)
    The model $\mathbf{f}$ is fitted by a multilayer perceptrong (MLP). The MSE is minimized using a version of the gradient descent algorithm (Adam, [49]) with weight decay. The MLP considered here always consisted of one hidden layer with 100 neurons. We found empirically that adding more layers can lead to instabilities even in stationary runs, while opting for a shallow, but wide network led to stable simulations. The activation function used is the swish function (SiLU), which we found yields smoother responses to external perturbations.
  2. ii.
    Causally constrained drift. Here we consider the same architecture of the unconstrained drift above. The main difference is in the loss function, here equal to:
    $\mathcal{L}=MSE(\Delta\mathbf{x},\mathbf{f}(\mathbf{x}))+\lambda\sum_{k,j\in\mathcal{S}}(\frac{\partial f_{k}}{\partial x_{j}})^{2}~,$ (10)
    where the set of entries $(k,j)$ is determined by the causal Adjacency matrix introduced in the main text. The derivative $\partial f_{k}/\partial x_{j}$ are computed with Autograd from PyTorch, providing automatic differentiation [81]. $\lambda$ is here set to $\lambda=10$.
Once a determistic drift $\mathbf{f}$ is computed we close the equation with a Wiener process. The first step is to compute the residuals of the deterministic dynamics as $\mathbf{r}(t)=\Delta\mathbf{x}-\mathbf{f}(\mathbf{x}(t))$, $\mathbf{r}(t)\in\mathbb{R}^{n}$. A stochastic parametrization for this residual term is then constructed. As the main paper is focused on proposing constraints acting on the drift term, we adopted the simplest option, in which the noise term in Eq. (8) is diagonal, with each entry $(i,i)$ given by the standard deviation of $r_{i}(t)$. For details on more complex stochastic closure, such as state-dependent noise we refer to the work of Falasca (2025) [29].
IV Charney De-Vore model: parameters
The values of the parameters used for the Charney-DeVore model (Eq. (3) in the main text) are reported in Table I below. All parameters are chosen with standard values as discussed in [23, 21, 26, 35]. The system is integrated using a Euler-Maruyama scheme with $dt=0.01$.
Table 1: Model coefficients for the six-dimensional stochastic Charney-DeVore model.
Parameter Value Description
$C$ 0.1 Newtonian relaxation rate
$x_{1}^{*}$ 0.95 Zonal background forcing (mode 1)
$x_{4}^{*}$ 0.76095-0.76095 Zonal background forcing (mode 4)
$\gamma$ $0.2$ Topographic height
$b$ $1.6$ Channel aspect ratio
$\alpha_{m}$ $\frac{8\sqrt{2}}{\pi}\frac{m^{2}}{4m^{2}-1}\frac{b^{2}+m^{2}-1}{b^{2}+m^{2}}$ Nonlinear advection (mode $m$)
$\beta_{m}$ $\frac{\beta b^{2}}{b^{2}+m^{2}}$ Coriolis effects (mode $m$)
$\delta_{m}$ $\frac{64\sqrt{2}}{15\pi}\frac{b^{2}+m^{2}+1}{b^{2}+m^{2}}$ Triad interaction (mode $m$)
$\gamma_{m}$ $\gamma\frac{4m^{3}}{4m^{2}-1}\frac{\sqrt{2}b}{\pi(b^{2}+m^{2})}$ Orographic damping (mode $m$)
$\tilde{\gamma}^{*}_{m}$ $\gamma\frac{4m}{4m^{2}-1}\frac{\sqrt{2}b}{\pi}$ Orographic forcing (mode $m$)
$\varepsilon$ $\frac{16\sqrt{2}}{5\pi}$ Wave-wave interaction
$\sigma$ 0.05 Noise amplitude
dtdt $0.01$ Integration time step
V Practical estimation of the response operator
To estimate the response operator in practice we follow the steps proposed in Appendix A of Falasca (2025) [29]. We report those steps here for completeness.
Given a model $d\mathbf{x}=\mathbf{f}(\mathbf{x})dt+\mathbf{\Sigma}d\mathbf{W}$, we build the impulse response operator as follows:
  1. 1.
    We simulate a very long trajectory of the model, starting from a random initial condition and remove an initial transient. We then sample $N_{e}$ random points from the simulation to define an ensemble of $N_{e}$ initial conditions on the model’s attractor. $N_{e}$ should be very large (i.e. $N_{e}\gg 1$) in order to sample the whole attractor and approximate averages over the invariant distribution. Here we chose $N_{e}=10^{6}$ members.
  2. 2.
    For each one of the $N_{e}$ initial conditions, we impose an impulse perturbation $\Delta_{j}$ to the degree of freedom $x_{j}$ at time $t=0$. The amplitude of the perturbation should be theoretically infinitesimally small, ensuring linearity of the response even in nonlinear systems. In practice we do as follows: we consider the long time series of $x_{j}(t)$ from the long control integration above and define $\Delta_{j}=10^{-2}\sigma_{j}$, where $\sigma_{j}$ is the standard deviation of time series $x_{j}(t)$.
  3. 3.
    Therefore, for a given initial condition $\mathbf{x}(0)$, we simulate two trajectories: a control trajectory without perturbation and a perturbed one, where an impulse perturbation $\Delta_{j}$ has been imposed on the $j$-th degree of freedom at time $t=0$. This procedure is repeated in parallel for all initial conditions, resulting in an ensemble of $N_{e}$ paired of control and perturbed trajectories.
  4. 4.
    At each time $t$ we then estimate the time dependent ensemble average $\langle A(x_{k}(t))\rangle$ of an observable $A(x_{k}(t))$ for both the perturbed and unperturbed run. We refer to the perturbed and control ensemble average respectively as $\langle A(x_{k}(t))\rangle_{\text{p}}$ and $\langle A(x_{k}(t))\rangle$. The observables considered in this study are going to be: (i) $A(x_{k}(t))=x_{k}(t)$ and (ii) $A(x_{k}(t))=(x_{k}(t)-\mu_{k}(t))^{2}$, with $\bm{\mu}(t)$ representing the time-dependent mean of the distribution, therefore quantifying the response in ensemble mean and variance.
  5. 5.
    We define the impulse response operator for observable $A(x_{k}(t))$ as
    $R_{k,j}(t)=\frac{\langle A(x_{k}(t))\rangle_{\text{p}}-\langle A(x_{k}(t))\rangle}{\Delta_{j}}.$
VI Neural emulators of the Charney-DeVore model: stationary statistics
Stationary density and autocorrelation functions of the Charney-deVore system as given by the numerical model, the vanilla emulator and the causal emulator.
Figure 3: Stationary statistics. First and second rows: stationary distributions (PDFs) of the variables $x_{1},...,x_{6}$ of the CdV model. Labels: “CdV” refers to the numerical model; “Model vanilla” is the vanilla, unconstrained emulator; “Model causal” is the causally constrained emulator. Third and fourth rows: same as the first row but for the autocorrelation functions (ACFs).
VII Response operator for the Charney-DeVore model
In Figure 4 and Figure 5, we report the estimation of the full response operator in ensemble mean and variance as obtained by the numerical model and the two emulators. The response operators have been computed with the method presented in Section V and using $N_{e}=10^{6}$ ensemble members. The unconstrained, vanilla emulator provides a very good representation of responses to impulse perturbations, as expected given the large training dataset and Markovianity of the system [29]. The causally constrained emulator yields a systematically improved representation of both mean and variance responses. This is further quantified by the total time-dependent MSE, aggregated over all $j$ and $k$, between the emulators and the ground truth. This analysis is shown in Figure 6.
Figure 4: Time-dependent, responses in ensemble mean of variable $x_{k}(t)$ to an impulse perturbation the CdV model in Eq. (3) in the main text as predicted by the unconstrained and causally constrained emulators. Example: column (1), row (2) represents what is the time-dependent response of $x_{k=2}(t)$ given a small, impulse perturbation imposed on $x_{j=1}(0)$ at time $t=0$. Labels: “CdV” refers to the numerical model; “Model vanilla” is the vanilla, unconstrained emulator; “Model causal” is the causally constrained emulator.
Figure 5: Time-dependent, responses in ensemble variance of variable $x_{k}(t)$ to an impulse perturbation the CdV model in Eq. (3) in the main text as predicted by the unconstrained and causally constrained emulators. Example: column (1), row (2) represents what is the time-dependent response of $x_{k=2}(t)$ given a small, impulse perturbation imposed on $x_{j=1}(0)$ at time $t=0$. Labels: “CdV” refers to the numerical model; “Model vanilla” is the vanilla, unconstrained emulator; “Model causal” is the causally constrained emulator.
Figure 6: Perturbed statistics in the linear regime. Panel (a): Total MSE over time for the mean response, computed from the full response operator $\mathbf{R}(t)\in\mathbb{R}^{6,6}$ for each model. Panel (b): Same as Panel (a) but the response in ensemble variance.
VIII Response to large step function forcings in the Charney-DeVore model
We now consider the nonlinear response regime by adding a step function forcing $\mathbf{F}$ to both the numerical (CdV) and neural models. In the main text we focused on the case $\dot{\mathbf{x}}=\mathbf{F}+\mathbf{f}(\mathbf{x})+\mathbf{\Sigma}\bm{\xi}(t)$, where $\mathbf{F}=(\sigma_{1},0,0,0,0,0)$ for $t\geq 0$, $\sigma_{1}$ being the std. dev. of the $x_{1}$ variable. We stress that this is a very large forcing, imposed on a model that had only access to unperturbed variability. Here, we perform six forcing experiments, applying a step forcing to each degree of freedom in turn with amplitude equal to its standard deviation: $\mathbf{F}=(\sigma_{1},0,\ldots,0)$, $(0,\sigma_{2},0,\ldots,0)$, $\ldots$, $(0,\ldots,0,\sigma_{6})$. For each case, we analyze the time-dependent response of the ensemble mean and variance of the full system. Results are shown in Fig. 7 (mean) and Fig. 8 (variance). The causally coarse-grained emulator outperforms the unconstrained, vanilla emulator. Notably, the vanilla emulator becomes unstable in three out of six forcing experiments (rows 1, 2, and 6 in Fig. 7), whereas the causally constrained emulator remains stable in all cases. Errors in the ensemble-mean response are primarily observed when the step forcing is applied to $x_{5}$ and $x_{6}$. To elucidate the origin of these errors, we revisit the response operator $R_{k,j}(t)$ inferred in the previous sections and focus on the ground-truth response estimated from the numerical model. We quantify the cumulative sensitivity of the system to perturbations applied to each degree of freedom $x_{j}$ as $D_{j}=\sum_{t=0}^{\tau_{\infty}}\sum_{k}R_{k,j}(t)$,; where is $\tau_{\infty}=40$ Model Time Units, consistent with all response analyses. The resulting sensitivities are shown in Fig. 9. This analysis identifies $x_{5}$ and $x_{6}$ as the most sensitive degrees of freedom: perturbations applied to these variables induce the largest cumulative system-wide response. This high sensitivity provides a plausible, qualitative explanation for the observed degradation in ensemble-mean responses under large forcings applied to $x_{5}$ and $x_{6}$, where even small modeling errors are strongly amplified.
Figure 7: Perturbed statistics in the nonlinear regime. Row (i) shows the repsonse in ensemble mean to a step function forcing applied on $x_{i}$; e.g. Row (1): response in ensemble mean response a step function forcing $\mathbf{F}=(\sigma_{1},0,0,0,0,0)$ applied for $t\geq 0$, where $\sigma_{1}$ is the standard deviation of $x_{1}$. Row (2): same as Row (1) but using $\mathbf{F}=(0,\sigma_{2},0,0,0,0)$. Labels: “CdV” refers to the numerical model; “Model vanilla” is the vanilla, unconstrained emulator; “Model causal” is the causally constrained emulator.
Figure 8: Perturbed statistics in the nonlinear regime. Row (i) shows the response in ensemble variance to a step function forcing applied on $x_{i}$; e.g. Row (1): response in ensemble mean to a step function forcing $\mathbf{F}=(\sigma_{1},0,0,0,0,0)$ applied for $t\geq 0$, where $\sigma_{1}$ is the standard deviation of $x_{1}$. Row (2): same as Row (1) but using $\mathbf{F}=(0,\sigma_{2},0,0,0,0)$. Responses are computed using an ensemble of $N_{e}=10^{5}$ members. Labels: “CdV” refers to the numerical model; “Model vanilla” is the vanilla, unconstrained emulator; “Model causal” is the causally constrained emulator.
Figure 9: Cumulative (causal) sensitivity $D_{j}=\sum_{t=0}^{\tau_{\infty}}\sum_{k}R_{k,j}(t)$ of the Charney-DeVore model in Eq. (3) in the main text. For the response operator $R_{k,j}(t)$ we use the correct ground-truth. The system is more sensitive to perturbations on degrees of freedom $x_{5}$ and $x_{6}$, therefore providing a possible explanation for the degradation of results when adding large, step function forcings on these degrees of freedoms.

### v2

@@
 These limitations are not restrictive for realistic turbulent system considered here, such as weather and climate, which lack exact symmetries, are only partially observed, and admit multiple effective coarse-grained representations.
-In this regime, dominant interactions among coarse-grained variables are often well approximated by a state-independent linear operator, with nonlinearities providing higher-order corrections [88, 87, 89, 90, 111, 41, 64, 70].
 There is no unique law to be discovered, instead the learned drift $\mathbf{f}$ represents one admissible effective realization, regularized to ensure physically consistent response behavior and agreement with observed statistics.
@@
 Accordingly, the causal constraints introduced in the main text, exploit vanishing entries of the ensemble-averaged Jacobian, $\langle\partial f_{k}(\mathbf{x})/\partial x_{j}\rangle=0$, to suppress the corresponding state-dependent couplings $\partial f_{k}(\mathbf{x})/\partial x_{j}$ in the neural network loss.
-In other words, the central assumption of this work is that knowledge of the zeros of the ensemble averaged Jacobian already provides an informative constraint for guiding the architecture of general neural emulators.
-A complication arises in idealized systems with exact symmetries, such as the Lorenz ’63 and Lorenz ’96 systems [58, 59].
+A complication can in principle arise in idealized systems with exact symmetries, such as the Lorenz ’63 and Lorenz ’96 systems [57, 58].
 In such cases $R_{k,j}(dt)=0$ can arise from exact cancellations in the ensemble average, even though $\partial f_{k}(\mathbf{x})/\partial x_{j}\neq 0$ pointwise.
-In such cases, a vanishing short-time response does not necessarily imply the absence of an instantaneous coupling.
-One possible strategy in such idealized settings is to explicitly reduce the symmetries of the underlying dynamical system, following symmetry-reduction approaches as advocated in [22].
-For example, in the Lorenz ’63 system [58], this corresponds to considering the so-called “proto-Lorenz” system [77], in which the symmetries of the original system are reduced.
-More importantly, however, we argue that this limitation does not constitute a practical obstacle for reduced-order modeling of real-world turbulent climate systems, which are the primary focus of this work.
-Such systems typically lack exact symmetries and admit no unique coarse-grained representation.
-The effective drift term $\mathbf{f}$ is validated on its ability to reproduce stationary and statistics given a specific choice of coarse-grained representation.
-In this context, we find it useful to provide additional qualitative justification for why focusing on zeros of the ensemble-averaged Jacobian is not a fundamental limitation.
-First, we discuss a qualitative mathematical argument in the context of linear inverse models of climate.
-Second, we emphasize a more general point regarding the non-uniqueness of effective dynamics in data-driven modeling of coarse-grained systems.
-  • In the context of reduced-order modeling of large-scale climate variability, typically constructed from anomalies, there is extensive evidence that linear inverse models (LIMs), i.e.
-$d\mathbf{x}=\mathbf{L}\mathbf{x}dt+\mathbf{\Sigma}d\mathbf{W}$, as originally proposed by Penland [88], provide a useful effective approximation of the dynamics and serve as powerful forecasting tools; see [88, 87, 89, 90, 111, 41] among many others.
-A natural generalization incorporates quadratic nonlinearities,
-    $d\mathbf{x}=(\mathbf{L}\mathbf{x}+\mathbf{B}(\mathbf{x},\mathbf{x}))dt+\mathbf{\Sigma}d\mathbf{W}$
-    as discussed, for example, in [64]. In what follows, we focus on nonlinear extensions of the LIM within this quadratic framework. In this case, the $k$-th component of the drift can be written as
-    $f_{k}(\mathbf{x})=\sum_{m}L_{k,m}x_{m}+\sum_{m,n}B_{k,m,n}x_{m}x_{n}~.$
-    and the corresponding Jacobian entry takes the form
-    $J_{k,j}(\mathbf{x})=\partial f_{k}(\mathbf{x})/\partial x_{j}=L_{k,j}+\sum_{m}B_{k,j,m}x_{m}+\sum_{n}B_{k,n,j}x_{n}~.$
-    Taking the ensemble average yields
-    $\langle J_{k,j}(\mathbf{x})\rangle=\langle\partial f_{k}(\mathbf{x})/\partial x_{j}\rangle=L_{k,j}+\sum_{m}B_{k,j,m}\langle x_{m}\rangle+\sum_{n}B_{k,n,j}\langle x_{n}\rangle~.$
-    In the common modeling regime where climate variability is represented in terms of anomalies, one typically has $\langle x_{n}\rangle=0$ for all components, leading to
-    $\langle J_{k,j}(\mathbf{x})\rangle=L_{k,j}~.$
-    In this modeling setting, the ensemble-averaged Jacobian therefore coincides with the linear operator $\mathbf{L}$.
-The empirical success of linear inverse models for coarse-grained climate variability suggests that, at this level of coarse-grained description (typically around 20 principal components of a given climate field, see e.g.
-[87]), the linear operator $\mathbf{L}$ approximates the dominant direct couplings, with nonlinear terms acting as corrections.
-This linear operator is obviously state independent.
-As a result, in this example, any vanishing entry $\langle J_{k,j}\rangle=L_{k,j}=0$ reflects the absence of a linear coupling, rather than a cancellation induced by ensemble averaging, therefore providing intuition for why vanishing entries of the ensemble-averaged Jacobian are structurally meaningful constraints.
-  • More broadly, this work highlights a principle relevant to data-driven modeling of realistic multiscale systems.
-As discussed in [29], the effective drift $\mathbf{f}$ inferred from data is inherently non-unique and depends on modeling choices such as variable selection, resolution, filtering, and projection onto latent representations.
-In this setting, there is no single “true” dynamics; instead, models are evaluated by their ability to reproduce stationary and perturbed statistics at the chosen level of coarse graining.
+One possible strategy in such idealized settings is to explicitly reduce the symmetries of the underlying dynamical system, following symmetry-reduction approaches advocated in [22].
+For example, Ref.
+[76] proposed a nonlinear coordinate transformation to quotient the $Z_{2}$ symmetry of the Lorenz ’63 system [57], leading to the so-called “proto-Lorenz” system.
+In the case of the Lorenz ’96 system, defined as $\dot{u_{j}}=(u_{j+1}-u_{j-2})u_{j-1}-u_{j}+F,~j=0,...,J$ with periodic boundary conditions, the model is translationally invariant, with each $u_{j}$ having the same ensemble mean $\langle u_{j}\rangle=\bar{u}$.
+In this latter case, introducing a site-dependent forcing $F_{j}$ is sufficient to guarantee the validity of our framework.
+For example, one could define $F_{j}=F+\eta_{j}$, with $\eta_{j}\sim\mathcal{N}(0,\sigma^{2})$, mimicking the effect of topography, and therefore leading to a more realistic forcing term.
+Crucially, we argue that the limitation discussed above does not constitute a practical obstacle for reduced-order modeling of realistic, complex turbulent dynamical systems, with no symmetries.
+As discussed in [29], in the case of climate modeling, the effective drift $\mathbf{f}$ inferred from data is inherently non-unique and depends on modeling choices such as variable selection, resolution, filtering, and projection onto latent representations.
+In this setting, there is no single “true” dynamics; instead, models are evaluated by their ability to reproduce stationary and perturbed statistics at the chosen level of coarse-graining.
 From this perspective, a vanishing short-time response $R_{k,j}(dt)=0$ provides sufficient justification to construct an effective model
@@
 (1) of the main text.
-Such comparisons should be performed at short time lags in order to avoid spurious responses that may arise at longer lags when using the FDT [28].
+Such comparisons should be performed at short time lags to avoid spurious responses that may arise at longer lags when using the FDT [28].
 The causal constraints introduced here should be applied when they improve the representation of the response operator relative to the unconstrained model.
-In systems with exact symmetries, as discussed above, any biases introduced by the constraints will already be apparent in the representation of stationary variability.
+In systems with exact symmetries, as discussed above, any biases introduced by the constraints will be immediately apparent in the representation of stationary variability.
 III Details on the emulator formulations

### v3

@@
-Causally constrained reduced-order neural models of complex turbulent dynamical systems
+Physics and causally constrained discrete-time neural models of turbulent dynamical systems
 Abstract
-We introduce a flexible framework based on response theory and score matching to suppress spurious, noncausal dependencies in reduced-order neural emulators of turbulent systems, focusing on climate dynamics as a proof-of-concept.
-We showcase the approach using the stochastic Charney–DeVore model as a relevant prototype for low-frequency atmospheric variability.
-We show that the resulting causal constraints enhance neural emulators’ ability to respond to both weak and strong external forcings, despite being trained exclusively on unforced data.
-The approach is broadly applicable to modeling complex turbulent dynamical systems in reduced spaces and can be readily integrated into general neural network architectures.
+We present a framework for constructing physics and causally constrained neural models of turbulent dynamical systems from data.
+We first formulate a finite-time flow map with strict energy-preserving nonlinearities for stable modeling of temporally discrete trajectories.
+We then impose causal constraints to suppress spurious interactions across degrees of freedom.
+The resulting neural models accurately capture stationary statistics and responses to both small and large external forcings.
+We demonstrate the framework on the stochastic Charney–DeVore equations and on a symmetry-broken Lorenz–96 system.
+The framework is broadly applicable to reduced-order modeling of turbulent dynamical systems from observational data.
 preprint: APS/123-QED
@@
 Modeling and understanding complex turbulent dynamical systems often requires abandoning full microscopic descriptions in favor of reduced-order models, where coarse-grained variables evolve according to effective stochastic dynamics [10, 59].
-Such reduced stochastic descriptions are fundamental across physics, ranging from chemical physics [47, 88], neuroscience [41, 89], mathematical physics [91, 90], turbulence [66, 94], climate science [42, 64, 69, 34, 59, 50, 49], and many others.
+Such reduced stochastic descriptions are fundamental across physics, ranging from chemical physics [41, 79], neuroscience [35, 80], mathematical physics [82, 81], turbulence [58, 84, 14], climate science [36, 57, 61, 28, 52, 43, 42], and many others.
 In practice, one is often faced with an inverse problem: given long trajectories of a Markovian, coarse-grained system, construct a data-driven model that faithfully reproduces its statistical and dynamical features [77, 37, 74, 65, 46, 44, 9, 88, 15, 14, 30, 2, 25, 23].
 Recent advances in machine learning further contribute to inverse modeling by providing flexible neural network architectures for data-driven emulation, with promising applications in fluid dynamics [27, 67, 94], weather [72, 6], and climate science [96, 95, 34, 12, 20].
-However, neural network–based models are typically agnostic to the underlying physics, which can lead to nonphysical behavior, such as violations of energy conservation - an issue that can be addressed through appropriate constraints [73, 51, 7, 77] - or incorrect responses to external forcings.
-The latter issue has been documented in neural weather models [101] and, especially, in climate emulators [102, 106].
-This limitation highlights a central open challenge for neural emulators of physical systems: high skill in reproducing stationary statistics does not guarantee accurate responses to external perturbations, thereby limiting the use of such models for process understanding through sensitivity analysis and the exploration of “what if?” scenarios.
-Crucially, a model’s ability to respond to forcings is inherently linked to causality, since cause–effect relationships in physical systems are inferred through external interventions [46, 2, 60, 62], commonly realized via perturbation–response experiments (see [82, 8] in the contexts of neuroscience and climate, respectively).
-This connection suggests that constraining a data-driven model to respect a given “correct” causal structure should improve its ability to respond to external forcings.
-The resulting challenge is twofold: to infer causal relationships from stationary, unperturbed data alone, and to incorporate this information into neural emulators, all within a computationally efficient framework that is both scalable and practical.
-In this context, the Fluctuation–Dissipation Theorem (FDT) plays a fundamental role as it provides a bridge between forced and natural fluctuations of physical systems [52, 30, 63, 74, 39, 61].
-The work of Baldovin et al.
-(2020) [5] showed that the FDT provides a rigorous causal inference strategy in the interventional sense directly from data, and this approach has since been adopted in a growing body of work [13, 4, 28, 27, 1, 36, 35, 29].
-In this Letter, we take this line of research one step further by leveraging the FDT to constrain the training of general neural emulators of physical systems.
-Given a trajectory $\mathbf{x}(t)=[x_{1},x_{2},...,x_{n}](t)$ of a multivariate Markovian system, our framework proceeds in two steps: (i) use the FDT to infer the direct causal couplings $x_{j}\rightarrow x_{k}$ across degrees of freedom $x_{j}$, $x_{k}$; (ii) learn a stochastic neural model from $\mathbf{x}(t)$ while preserving the inferred causal structure.
-A related direction was recently proposed by Chen and Liu (2025) [16], who used causation entropy to rank candidate functions in the construction of causally constrained reduced-order models, highlighting the benefits of causally informed approaches.
-Our work addresses a complementary setting by providing a practical strategy to enforce a prescribed causal structure within general neural network architectures.
-To illustrate the methodology, we focus on climate dynamics as a representative example of a complex, turbulent dynamical system that is necessarily modeled through an effective, coarse-grained description [69, 51, 66], and for which accurate modeling of forced responses is essential [29].
-We consider a stochastic version of the Charney–DeVore model [15], relevant for conceptual studies of low-frequency climate variability.
-This choice provides a simplified testbed that nevertheless retains key features of realistic coarse-grained climate dynamics, including strong mixing, fast decay of autocorrelations, and regular, unimodal distributions [70, 68].
-We show that causally constrained neural emulators substantially improve the ability of neural networks to capture responses to both weak and strong forcings.
-In particular, the causal emulator accurately reproduces responses beyond the linear regime, including time-dependent shifts of the attractor in state space, despite being trained solely on stationary, unperturbed data.
-II Framework: from linear response to causal constraints
-Fluctuation–Dissipation Theorem (FDT) and causality
-Given an nn–dimensional stochastic nonlinear system $\mathbf{x}(t)$, the FDT [63, 74] states that the linear response, in the ensemble average sense, of a degree of freedom $x_{k}(t)$ to an infinitesimally small impulse perturbation $\delta x_{j}(0)$ applied to component $x_{j}(0)$ can be written as:
-$\begin{split}R_{k,j}(t)&=\frac{\delta\langle x_{k}(t)\rangle}{\delta x_{j}(0)}=-\langle x_{k}(t)s_{j}(\mathbf{x}(0))\rangle,\\ s(\mathbf{x})&=\nabla\log\rho(\mathbf{x}).\end{split}$ (1)
+However, neural network–based models are typically agnostic to the underlying physics.
+This can lead to several failure modes, including long-term nonphysical behavior, such as violations of energy conservation [65, 44, 69] and consequent blow-up of solutions [75], and incorrect responses to external forcings [90, 93, 97].
+The latter limitation is particularly critical for scientific applications, where models are used for process understanding and exploration of “what if?” scenarios.
+Crucially, a model’s ability to respond to forcings is inherently linked to causality, since cause–effect relationships in physical systems are defined by their response to external interventions [40, 1, 53, 55], typically realized via perturbation–response experiments [73, 5].
+In this Letter, we address these challenges by presenting an efficient strategy to construct physics and causally constrained neural models of turbulent dynamical systems from data alone.
+We focus on the less explored context of discrete-time modeling, which avoids inconsistencies arising in applications when continuous-time conservation principles are applied to temporally coarse-grained data.
+Given a trajectory $\mathbf{x}_{t}=(x^{(1)}_{t},x^{(2)}_{t},...,x^{(n)}_{t})$ of a multivariate discrete Markovian system, our framework proceeds in two steps: (i) Physics Constraints: We propose a neural model, with strict energy-conserving nonlinearities, tailored to realistic, discrete-time coarse-grained observations.
+(ii) Causal Constraints: We identify direct causal couplings $x^{(j)}\rightarrow x^{(k)}$ from data via the Fluctuation-Dissipation Theorem (FDT) and use this information to suppress spurious, non-causal interactions during model training.
+The FDT plays a fundamental role here, as it provides a rigorous bridge between forced and natural fluctuations [45, 56, 54], allowing for a causal inference strategy in the interventional sense directly from unperturbed, stationary data [3, 11, 23].
+Ultimately, the physics constraints, specifically formulated for discrete settings, provide the foundation of our approach.
+These physical priors promote robust inference in data-scarce regimes and ensure long-term numerical stability.
+Furthermore, this structural rigidity naturally bounds the impact of imperfect causal estimates: even if the inferred causal links contain errors, the resulting model remains confined to a physically consistent solution.
+As a representative application, we consider climate dynamics, where accurate prediction of forced responses is essential [23].
+We test the framework on two benchmarks: a stochastic version of the Charney–DeVore model [13] and a symmetry-broken Lorenz–96 system [51] as a higher-dimensional testbed.
+The resulting neural emulators are stable, capture stationary statistics, and reproduce responses to both weak and strong external forcings, including time-dependent shifts of the attractor, despite being trained exclusively on unperturbed data.
+II Framework: physics and causal constraints
+II.1 Finite-time flow map with energy-preserving nonlinearities
+Previous work.
+We build on the abstract representation of turbulent dynamical systems often adopted by Majda and collaborators [56, 59] $\dot{\mathbf{x}}=\mathbf{F}+\mathbf{A}\mathbf{x}+\mathbf{B}(\mathbf{x},\mathbf{x})+\mathbf{\Sigma}\bm{\xi}(t)$, which arises when a large class of fluid flows are projected onto orthogonal basis functions.
+The quadratic term is energy-conserving, satisfying $\mathbf{x}\cdot\mathbf{B}(\mathbf{x},\mathbf{x})=0$ [50, 47], with the system’s energy defined as $E=(1/2)\mathbf{x}^{\mathrm{T}}\mathbf{x}$.
+$\mathbf{F}$ represents external forcing, and the linear operator $\mathbf{A}$ is typically decomposed as $\mathbf{A}=\mathbf{L}+\mathbf{D}$, where $\mathbf{L}$ (skew-symmetric) and $\mathbf{D}$ (symmetric, negative-definite) represent dispersion and dissipation, respectively.
+$\mathbf{\Sigma}\bm{\xi}(t)$ is a stochastic forcing, with $\bm{\xi}(t)$ denoting a standard Gaussian white-noise process.
+Following recent work on neural closure and modeling [92, 84, 83, 30], we replace $\mathbf{B}(\mathbf{x},\mathbf{x})$ by a general nonlinear operator $\mathbf{S}(\mathbf{x})\mathbf{x}$, where $\mathbf{S}(\mathbf{x})$ is constrained to be skew-symmetric, $\mathbf{S}^{\mathrm{T}}(\mathbf{x})=-\mathbf{S}(\mathbf{x})$.
+This preserves the key physical property $\mathbf{x}^{\mathrm{T}}\mathbf{S}(\mathbf{x})\mathbf{x}=0$, so the nonlinear term redistributes energy across modes without net injection or dissipation.
+The importance of energy-conserving nonlinearities in data-driven models of turbulent systems was highlighted in [65].
+Details are presented in the Supplemental Material (SM) Sec.
+I.
+Finite-time flow map formulation.
+Observational data are fundamentally discrete and temporally coarse-grained, yet physics constraints are typically defined strictly in the continuous limit.
+To resolve this discrepancy, we introduce a discrete, abstract formulation of turbulent dynamical systems with exact energy-conserving nonlinearities for arbitrary time increments:
+$\mathbf{x}_{t+1}=\mathbf{f}(\mathbf{x}_{t})=\mathbf{F}+\mathbf{M}\mathbf{Q}(\mathbf{x}_{t})\mathbf{x}_{t}+\mathbf{\Sigma}\bm{\xi}_{t},$ (1)
+where $\mathbf{F}$ and $\mathbf{\Sigma}\bm{\xi}_{t}$ represent the effective deterministic and stochastic forcings integrated over the finite time step.
+The nonlinear operator $\mathbf{Q}(\mathbf{x}_{t})$ is constrained to be strictly orthogonal ($\mathbf{Q}^{\top}\mathbf{Q}=\mathbf{I}$).
+Thus, the deterministic map splits into an energy-preserving rotation, $\mathbf{v}_{t}=\mathbf{Q}(\mathbf{x}_{t})\mathbf{x}_{t}$, followed by a linear transformation $\mathbf{M}\mathbf{v}_{t}$.
+The isolated nonlinear term contributes no net growth or decay to the energy budget, strictly preserving the $L_{2}$-norm of the state:
+$\|\mathbf{Q}(\mathbf{x}_{t})\mathbf{x}_{t}\|^{2}=\mathbf{x}_{t}^{\top}\mathbf{Q}^{\top}(\mathbf{x}_{t})\mathbf{Q}(\mathbf{x}_{t})\mathbf{x}_{t}=\|\mathbf{x}_{t}\|^{2}.$
+Consequently, energy growth in the forward map is bounded by the dominant singular value of $\mathbf{M}$.
+While additional constraints on $\mathbf{M}$ could be imposed, they are often not required in practice and may degrade performance.
+In the SM (Sec.
+II), we provide mathematical justification for Eq.
+(1).
+First, we demonstrate that in the continuous limit this mapping arises as a Lie–Trotter splitting scheme [7], and we further derive the corresponding effective deterministic and stochastic forcings $\mathbf{F}$ and $\mathbf{\Sigma}\bm{\xi}_{t}$.
+As illustrative examples, we recast the deterministic Lorenz-63 system and a stochastic triad model within this framework.
+Second, we return to the data-driven setting and show that the proposed discrete formulation in Eq.
+(1) enables stable learning from severely sub-sampled data, where both numerical integrators and continuous-time data-driven models fail.
+Neural model fitting.
+Given a stationary trajectory $\mathbf{x}_{t}$, we parametrize the deterministic drift $\mathbf{f}(\mathbf{x}_{t})$ by training a multilayer perceptron (MLP) to minimize the mean squared error $\text{MSE}(\mathbf{x}_{t+1},\mathbf{f}(\mathbf{x}_{t}))$.
+States are standardized to zero mean and unit variance.
+We enforce the orthogonality of $\mathbf{Q}(\mathbf{x}_{t})$ by predicting the matrix exponential $\mathbf{Q}(\mathbf{x}_{t})=\exp(\mathbf{S}(\mathbf{x}_{t}))$ of a skew-symmetric matrix $\mathbf{S}(\mathbf{x}_{t})$.
+Crucially, the linear terms are initialized with a “first guess” ordinary least squares (OLS) solution, while the MLP is initialized to yield $\mathbf{Q}(\mathbf{x}_{t})=\mathbf{I}$.
+This ensures the joint optimization starts from a Linear Inverse Model (LIM) [77] baseline before learning the nonlinear corrections.
+Here, the diffusion matrix $\mathbf{\Sigma}$ is estimated as a diagonal matrix from the standard deviation of the determistic residuals; the general framework captures spatial correlations via the Cholesky decomposition of the residual covariance.
+Further implementation details, including careful data rescaling, are discussed in SM Sec.
+I.
+II.2 Fluctuation–Dissipation Theorem (FDT) and causality
+Given an nn–dimensional stochastic nonlinear system $\mathbf{x}_{t}=(x^{(1)}_{t},x^{(2)}_{t},...,x^{(n)}_{t})$, the FDT [56, 66] states that the linear response, in the ensemble average sense, of a degree of freedom $x^{(k)}_{t}$ to an infinitesimally small impulse perturbation $\delta x^{(j)}_{0}$ applied to component $x^{(j)}_{0}$ can be written as:
+$\begin{split}R^{k,j}_{t}&=\frac{\delta\langle x^{(k)}_{t}\rangle}{\delta x^{(j)}_{0}}=-\langle x^{(k)}_{t}s_{j}(\mathbf{x}_{0})\rangle,\\ s(\mathbf{x})&=\nabla\log\rho(\mathbf{x}).\end{split}$ (2)
 $\rho(\mathbf{x})$ is the system’s invariant density and $s(\mathbf{x})$ is the score function of the system [31, 29].
-The focus on stochastic systems allows us to circumvent difficulties with determistic dissipative dynamical systems, where the invariant measure is singular [93, 92, 19].
 The brackets $\langle\cdot\rangle$ represents ensemble averages, computed over the invariant density.
 Eq.
-(1) provides a rigorous strategy for inferring causal relations in physical systems in the interventional sense from data alone [5].
-As discussed in [29], $R_{k,j}(t)$ supports multiple causal interpretations: values at the shortest time scale ($t=dt$) encode direct causal links $x_{j}\rightarrow x_{k}$.
-Nonzero responses at later times reflect indirect causal links.
-Additionally, a cumulative measure of causal influence can be defined as $D_{k,j}=\int_{0}^{\infty}R_{k,j}(t)dt$ [5].
-The causal constraints proposed in this work focus on the direct causal links, $R_{k,j}(dt)$.
-In practice the response is computed for standardized data as $(\sigma_{j}/\sigma_{k})R_{k,j}(dt)$, $\sigma_{i}$ representing the standard deviation of $x_{i}$; for simplicity, we denote this standardized operator by $R_{k,j}$.
-FDT from data
-The main challenge in estimating responses via Eq.
-(1) lies in estimating the score function $s(\mathbf{x})$, particularly in high-dimensional settings.
-Traditional studies have adopted the quasi-Gaussian (qG, hereafter) ansatz [55, 63], in which $\rho$ is approximated by a Gaussian.
-The qG approximation has proven effective for predicting mean responses, even for nonlinear problems [63, 39, 70, 35].
-Recent developments in score-based generative modeling [96, 95], offer new ways forward to overcome previous approximations by parametrizing the score function directly from data.
-Importantly, when provided with enough data, these new strategies scale to high-dimensional systems such as quasi-geostrophic flows [84], idelized ocean flows [97], weather [87], climate [10], fluid dynamics [56], among others.
-The estimated score $s(\mathbf{x})$ can then be inserted into the analytical response formula in Eq.
-(1), circumventing the need for explicit approximations of the invariant density, as in the quasi-Gaussian case.
-This approach was first proposed in [37].
-These advances are leading to a set of score-estimation techniques rather than a single strategy.
-For example, [35] highlighted different denoising score-matching procedures tailored to the dimensionality of the underlying system.
-Here, we further contribute to this “FDT–generative modeling toolbox”.
-Specifically, we adopt the score-matching estimator proposed in [45], allowing us to infer the score directly on the attractor, rather than up to a small noise perturbation (e.g., as in denoising score matching).
-The estimated score is used as an initial approximation and is then refined using the correction procedure introduced in [37].
-Details are reported in Section I of the Supplemental Material.
+(2) provides a rigorous strategy for inferring causal relations in physical systems in the interventional sense from data alone [3, 23].
+Values at the shortest time scale $R^{k,j}_{1}$ encode direct causal links $x^{(j)}\rightarrow x^{(k)}$ and are used to construct our causal constraints.
+Evaluating Eq.
+(2) requires the score function $s(\mathbf{x})$.
+We leverage recent advances in score-based generative modeling [86, 85, 75, 87, 78, 8, 49] to learn $s(\mathbf{x})$ directly from data for use in the analytical FDT formula [31, 29], therefore bypassing traditional quasi-Gaussian approximations [48, 56].
+Specifically, we employ the exact score-matching estimator [39] to evaluate the score directly on the system’s attractor.
+The resulting estimate is further refined using the correction procedure of [31] (see SM Sec.
+III for details).
 Causal constraints.
-Consider a stochastic dynamical system
-$\dot{x_{i}}(t)=f_{i}(x_{1},x_{2},...,x_{n})+\sigma_{i}\xi_{i}(t),~\text{with}~i=1,2,...,n,$
-where $\sigma_{i}\xi_{i}(t)$ is a stochastic forcing, $\bm{\xi}(t)$ being a standard Gaussian white noise process.
-A perturbation to $x_{j}$ can affect $x_{k}$ at the next time step only if $f_{k}$ depends on $x_{j}$.
-Thus, in the limit $dt\to 0$, $\partial f_{k}/\partial x_{j}\equiv 0$ implies the absence of a direct causal link and consequently, a vanishing short-time response $R_{k,j}(dt)=0$.
+Given a discrete stochastic systems $x^{(i)}_{t+1}=f_{i}(x^{(1)}_{t},x^{(2)}_{t},...,x^{(n)}_{t})+\sigma_{i}\xi^{(i)}_{t}~\text{with}~i=1,2,...,n$, a perturbation to $x^{(j)}$ can affect $x^{(k)}$ at the next time step only if $f_{k}$ depends on $x^{(j)}$.
+Thus, $\partial f_{k}/\partial x^{(j)}\equiv 0$ implies the absence of a direct causal link and consequently, a vanishing short-time response $R^{k,j}_{1}=0$.
 In practice, the system’s Jacobian is not observable from data, and a reliable numerical estimation is unfeasible in high-dimensional systems.
-However, the short-time response $R_{k,j}(dt)$ can be efficiently estimated from unperturbed data using the FDT in Eq.
-(1) as shown in [35].
-A vanishing response at the shortest resolved time scale thus implies $\big\langle\partial f_{k}(\mathbf{x})/\partial x_{j}\big\rangle=0$ for $j\neq k$.
-This information on vanishing entries of the ensemble-averaged Jacobian is used to constrain neural architectures.
-Thus, given a long trajectory of $\mathbf{x}(t)$, we first estimate the response operator $\mathbf{R}(dt)\in\mathbb{R}^{n,n}$ with the score-matching procedure introduced earlier.
-We then proceed by constructing a binary adjacency matrix $\mathbf{A}\in\mathbb{R}^{n,n}$, by thresholding $\mathbf{R}(dt)$.
-$A_{k,j}=0$ denoting the absence of a direct causal link $x_{j}\rightarrow x_{k}$, and $A_{k,j}=1$ denoting its presence.
-One way to threshold $R_{k,j}(dt)$ is to examine its values on a logarithmic scale; a visual inspection of the response magnitudes often reveals a clear separation between significant links and noise.
-The absence of such separation may itself be informative, suggesting a densely connected system, in which case no additional causal constraints are needed.
+However, the short-time response $R^{k,j}_{1}$ can be efficiently estimated from unperturbed data using the FDT in Eq.
+(2).
+A vanishing response at the shortest resolved time scale thus implies $\big\langle\partial f_{k}(\mathbf{x})/\partial x^{(j)}\big\rangle=0$ for $j\neq k$.
+This information is used to constrain neural architectures.
+Thus, given a long trajectory of $\mathbf{x}(t)$, we first estimate the short-time response operator $\mathbf{R}_{t=1}\in\mathbb{R}^{n,n}$.
+We then construct a binary adjacency matrix $\mathbf{A}\in\mathbb{R}^{n,n}$, by thresholding $\mathbf{R}_{1}$.
+$A^{k,j}=0$ denoting the absence of a direct causal link $x^{(j)}\rightarrow x^{(k)}$, and $A^{k,j}=1$ denoting its presence.
+A visual inspection of $R^{k,j}_{1}$ on a logarithmic scale often reveals a clear separation between significant links and noise.
+The absence of such separation suggests a densely connected system with no need for causal constraints.
 To provide an unsupervised criterion, we propose the following procedure:
   1. i.
-    Self-interactions. We assume that for any dissipative dynamical system $x_{j}\rightarrow x_{j}$ and therefore $A_{j,j}=1$.
+    Self-interactions. We assume that for any dissipative dynamical system $x^{(j)}\rightarrow x^{(k)}$ and therefore $A^{j,j}=1$.
   2. ii.
     Cross-interactions.
-We consider the off-diagonal values of $R_{k,j}(dt),j\neq k$, and apply a logarithmic transformation $\ln|R_{k,j}(dt)|$.
-This helps separate potential causal links (larger responses) from the bulk of spurious, near-zero interactions.
-A $k$-means algorithm [85] ($k=2$) is then used to partition these log-scaled values into two clusters, corresponding to “significant” and “insignificant” responses.
-An entry of $A_{k,j}=1$ is set if its corresponding $\ln|R_{k,j}(dt)|$ value belongs to the cluster with the larger centroid; otherwise, it is set to 0.
-The causal Adjacency matrix is then used to constrain the training of a neural emulator.
-Given a $n$-dimensional trajectory $\mathbf{x}(t)=[x_{1},x_{2},...,x_{n}](t)$ we fit a neural operator $\mathbf{f}$ such that $\dot{\mathbf{x}}(t)=\mathbf{f}(\mathbf{x})+\mathbf{\Sigma}\bm{\xi}(t)$.
-The causal constraints are imposed on the deterministic drift $\mathbf{f}$.
-We learn $\mathbf{f}$ via a mean squared error (MSE) minimization such that $\Delta\mathbf{x}=\mathbf{x}(t+1)-\mathbf{x}(t)=\mathbf{f}(\mathbf{x}(t))$, where the time unit has been rescaled by the time sampling interval.
-Details on the learning process are found in the SM, Section III.
-A “vanilla”, unconstrained operator $\mathbf{f}$ is trained by minimizing MSE as $\mathcal{L}=MSE(\Delta\mathbf{x},\mathbf{f}(\mathbf{x}))$.
-The causally constrained emulator is trained by minimizing the following loss function:
-$\mathcal{L}=MSE(\Delta\mathbf{x},\mathbf{f}(\mathbf{x}(t)))+\lambda\sum_{k,j\in\mathcal{S}}(\frac{\partial f_{k}}{\partial x_{j}})^{2}~,$ (2)
-where $\mathcal{S}$ is the set of pairs $(k,j)$ for which the causal adjacency matrix $A_{k,j}=0$.
-$\lambda$ is a parameter determining the strength of the regularization.
-The derivatives are obtained via automatic differentiation [80].
-We adopt soft constraints for two main reasons.
-First, they are computationally efficient [77].
-Second, in real-world, high-dimensional climate systems with partial observations and finite data, causal estimates are inherently uncertain.
-In this regime, hard constraints based on imperfect causal information will introduce biases into the learned dynamics.
-Soft constraints, in contrast, provide a flexible form of regularization, ensuring an appropriate balance between data fidelity, as captured by the MSE, and causal regularization.
-Additionally, we note that the regularization in Eq.
-(10) can be adapted to novel architectures, including generative models [18], which we leave to future work.
+We consider the off-diagonal values of $R^{k,j}_{1},j\neq k$, and apply a logarithmic transformation $\ln|R^{k,j}_{1}|$, separating potential causal links (larger responses) from the bulk of spurious, near-zero interactions.
+A k-means algorithm [76] (k=2) is used to partition these values into two clusters.
+An entry of $A^{k,j}=1$ is set if its corresponding $\ln|R^{k,j}_{1}|$ value belongs to the cluster with the larger centroid; otherwise, it is set to 0 (absence of causal link $x^{(j)}\rightarrow x^{(k)}$).
+The matrix $A^{k,j}$ is then used to constrain neural emulators. Specifically, the MSE minimization is regularized with a causal penalty:
+$\mathcal{L}$ $=\mathcal{L}_{\text{MSE}}+\mathcal{L}_{Causal}$ (3)
+$=\text{MSE}(\mathbf{x}_{t+1},\mathbf{f}(\mathbf{x}_{t}))+\lambda\sum_{k,j\in\mathcal{S}}(\frac{\partial f_{k}}{\partial x^{(j)}})^{2}~,$
+where $\mathcal{S}$ is the set of pairs $(k,j)$ for which $A_{k,j}=0$. $\lambda$ is a parameter determining the strength of the regularization.
+Imperfect causal estimates.
+False negatives in the adjacency matrix $A^{k,j}$ (true links penalized as noncausal) can degrade the learned emulator when learning via Eq.
+(3).
+Physics-constrained architectures mitigates the impact of imperfect causal estimates through rigid structural priors.
+Additionally, if false negatives are suspected, via validation with stationary and perturbed statistics (see paragraph below), a more robust penalty can be employed:
+$\mathcal{L}_{Causal}^{Robust}=\lambda\sum_{k,j\in\mathcal{S}}\min\!\Big((\frac{\partial f_{k}}{\partial x^{(j)}})^{2},\gamma\Big),$ (4)
+which bounds the penalty once $(\partial f_{k}/\partial x^{(j)})^{2}\geq\gamma$, allowing the MSE term to dominate when interactions are mistakenly excluded.
+An heuristic estimate for $\gamma$ based on the FDT operator $R^{k,j}_{1}$ is given in Sec.
+IC in the SM.
 Possible limitations for idealized systems.
-The proposed constraints leverage information about zeros of the ensemble-averaged Jacobian to constrain the state-dependent Jacobian.
-While this choice is practical and efficient, it can be limiting in idealized systems, for example because of exact symmetries.
-In such cases, nonzero local couplings may cancel in ensemble averages, yielding $\langle\partial f_{k}(\mathbf{x})/\partial x_{j}\rangle=0$ even if $\partial f_{k}(\mathbf{x})/\partial x_{j}\neq 0$, as occurs, for example, in the Lorenz ’63 model [57].
-In such idealized settings, applying the proposed approach may require a symmetry-reduction step, as advocated in [22] and demonstrated in [76] for the Lorenz ’63 system.
-These limitations are not restrictive for realistic turbulent system considered here, such as weather and climate, which lack exact symmetries, are only partially observed, and admit multiple effective coarse-grained representations.
-There is no unique law to be discovered, instead the learned drift $\mathbf{f}$ represents one admissible effective realization, regularized to ensure physically consistent response behavior and agreement with observed statistics.
-Further details are provided in Section II of the SM.
-Practical assessment of causal constraints from data alone.
-Finally, the utility of causal constraints for a given system can be readily assessed by fitting a constrained model and an unconstrained counterpart.
-A key diagnostic is the comparison of the response operator computed from the emulators (see Section V of the SM) with the FDT-based operator in Eq.
-(1).
-When the inferred causal structure is informative, the constrained emulator improves the representation of linear responses without degrading the representation of stationary statistics.
-Moreover, comparing modeled and FDT-inferred response operators provides a natural criterion for a posteriori calibration of neural emulators based on forced responses.
-Related response-based calibration ideas have been explored in previous studies [71, 72]; a systematic formulation tailored to neural emulators and causal constraints is left for future work.
-III Causal constraints in practice
-Test-model for low-frequency climate variability.
-To showcase the framework, we consider a stochastic version of the Charney–DeVore model [15], “CdV” hereafter.
+Causal constraints rely on zeros of the ensemble-averaged Jacobian to restrict the state-dependent Jacobian.
+In idealized systems with exact symmetries (e.g., the Lorenz ’63 system [50]), nonzero local couplings may cancel in the average, yielding $\langle\partial f_{k}(\mathbf{x})/\partial x^{(j)}\rangle=0$ even when $\partial f_{k}/\partial x^{(j)}\neq 0$.
+Such cases require a symmetry-reduction step [18, 68].
+However, this is rarely restrictive for realistic, partially observed turbulent flows (e.g., weather and climate), which lack exact symmetries and admit multiple coarse-grained representations, see Sec.
+IV in the SM for details.
+Assessing causal constraints.
+Unlike rigid physical priors, the utility of causal regularization is system-dependent and contingent on the accuracy of the data-driven causal discovery.
+The benefit of augmenting physics-constrained models with causal information can be assessed directly from data by comparing the FDT-based operator (Eq.
+(2)) with the emulators’ response operators (SM Sec.
+V).
+When the inferred causal structure is informative, the resulting emulator improves linear response predictions without degrading stationary statistics.
+III Physics and causal constraints in practice
+We first illustrate the proposed framework on a six-dimensional stochastic system, highlighting the causal inference procedure and the benefits of the imposed constraints.
+We then focus on a symmetry-broken Lorenz–96 system as a more demanding test -case, evaluating both its response to strong external forcings and robustness against imperfect causal estimates.
+III.1 Stochastic Charney-DeVore model
+We consider a stochastic version of the Charney–DeVore model [13], “CdV” hereafter, later reformulated by De Swart [19].
 The model was obtained as a six-dimensional Galerkin truncation of the barotropic vorticity equation on a $\beta$-plane channel with topography.
-The formulation considered here is the one further proposed by De Swart [23].
-The model has been widely studied in the context of large-scale atmospheric circulation [20, 21, 26], as well as in computational studies of regime transitions and rare events [38].
-The model takes the following form:
+The model has been studied in the context of large-scale atmospheric circulation [16, 17, 21, 32] and takes the following form:
 $\dot{x}_{1}$ $=\tilde{\gamma}_{1}x_{3}-C(x_{1}-x_{1}^{*})+\sigma\xi_{1}(t),$ (5)
@@
 $\dot{x}_{6}$ $=(\alpha_{2}x_{1}-\beta_{2})x_{5}-\gamma_{2}x_{4}-Cx_{6}+\delta_{2}x_{4}x_{2}+\sigma\xi_{6}(t).$
-In what follows, the six variables $x_{i}$ serve as an effective reduced-order description of an underlying high-dimensional dynamical system, from which we aim to learn a causally constrained neural emulator.
-The parameters and their physical interpretation are reported in Section IV of the SM.
-$\xi_{i}(t),i=1,...,6$ is a Gaussian white noise with zero mean, unit variance, and $\sigma$ is the noise amplitude.
-This stochastic forcing mimics the effect of unresolved fast processes.
-The model is an ideal first test case for our framework of large-scale, low-frequency climate dynamics for two main reasons.
-First, Eq.
-(3) features the abstract form $\dot{\mathbf{x}}=\mathbf{F}+\mathbf{L}\mathbf{x}+\mathbf{B}(\mathbf{x},\mathbf{x})+\mathbf{\Sigma}\bm{\xi}(t)$, with energy conserving quadratic nonlinearities such that $\mathbf{x}\cdot\mathbf{B}(\mathbf{x},\mathbf{x})=0$; a structure that naturally arises when geophysical turbulent flows are projected onto orthogonal bases [63, 64, 65, 66].
-Second, this system is strongly mixing, with fast decay of autocorrelations and regular, unimodal probability distributions: such properties are relevant for coarse-grained climate dynamics as found both in models and observational data [32, 6, 100, 68], and underlie the reduced-order modeling of realistic, complex turbulent systems [67].
-In this respect, the resulting dynamics differ qualitatively from those of prototypical chaotic models like the Lorenz ’63 system [57], which is weakly mixing, possesses exact symmetries, and exhibits a singular invariant measure [68].
-We simulate this model for $T=5\cdot 10^{7}$ time-steps at $dt=0.01$, i.e.
-$5\cdot 10^{5}$ Model Time Units (MTU).
-This long integration serves as a stationary run to (i) infer the response operators $R_{k,j}(t)$ from data alone and (ii) to build neural emulators.
-Two emulators, with and without causal constraints, are then compared in terms of their ability to reproduce both stationary and perturbed statistics.
-While a comprehensive analysis is provided in the Supplemental Material, the main text highlights the identification of causal links and the ability of the causally constrained emulator to capture time-dependent responses in ensemble mean and variance under strong forcings.
+We integrate Eq.
+(5) with a RK4 scheme for $T=10^{7}$ time-steps at $dt=0.01$, i.e.
+$10^{5}$ Model Time Units (MTU).
+The resulting 6-dimensional trajectory is then treated as a discrete time series $\mathbf{x}_{t}=(x^{(1)}_{t},...,x^{(6)}_{t})$ from which we want to learn the discrete flow map in Eq.
+(1).
+The parameters and their physical interpretation are reported in Sec.
+V of the SM.
+This stochastic model features fast decay of autocorrelations and regular, unimodal probability distributions: such properties are relevant for coarse-grained climate dynamics as found both in models and observations [26, 4, 89, 60], and underlie the reduced-order modeling of realistic, turbulent systems [59].
+The main text highlights the identification of the causal Adjacency matrix and response to weak and strong forcing of variable $x_{1}$; a comprehensive analysis is provided in the SM.
 Causal Adjacency matrix from data.
-Given a long trajectory $\mathbf{x}(t)=[x_{1},...,x_{6}](t)$ of the CdV model, we infer the direct causal links through the steps proposed in the previous Section.
-The strategy is highlighted step by step in Figure 1.
-We infer the response operator at the shortest time scale $R_{k,j}(dt)=-\langle x_{k}(dt)s_{j}(\mathbf{x}(0))\rangle$ via score-matching, consider only the off-diagonal elements, and apply the transformation $\ln|R_{k,j}(dt)|$; see Figure 1(a).
-We then apply a $k$-means algorithm ($k=2$) to split the large responses from the spurious non-causal ones; see Figure 1(b).
-This information is used to define the Adjacency matrix in Figure 1(c), thus defining the causal functional form of the system to be imposed in the neural emulator; see 1(d).
-In Section VII of the SM we also report results for the inference of the response operator $R_{k,j}(t)$ over multiple time lags $t$ using both the score-matching and quasi-Gaussian approximations, further illustrating the advantages of the score-matching approach.
+Given a trajectory of the CdV model, we infer the direct causal links.
+The strategy is summarized in Figure 1.
+We infer the response operator at the shortest time scale $R^{k,j}_{1}=-\langle x^{(k)}_{1}s_{j}(\mathbf{x}_{0})\rangle$, apply the transformation $\ln|R^{k,j}_{1}|$ and split the large responses from the spurious non-causal ones via k-means clustering (k=2); see Figure 1(a).
+This information is used to define the Adjacency matrix in Figure 1(b).
 Figure 1: Causal constraining of the CdV model (Eq.
-(3)) from data through the FDT.
-Panel (a): The 30 off-diagonal responses $\ln|R_{k,j}(dt)|$ (evaluated at the shortest time scale $t=dt$), flattened and plotted against a dummy index.
-Panel (b): Unsupervised identification of a response threshold via $k$-means clustering ($k=2$): clusters are colored in black and blue.
-Panel (c): Causal adjacency matrix: $A_{k,j}=1$ is assigned if its corresponding response value lies in the higher centroid cluster; otherwise $A_{k,j}=0$.
-Panel (d): “causal” deterministic functional form to be enforced in the emulator loss function in Eq.
+(5)) from data using the FDT and score matching.
+Panel (a): The 30 off-diagonal responses $\ln|R^{k,j}_{1}|$, flattened and plotted against a dummy index.
+Causal links are identified via k-means clustering ($k=2$), shown as filled and open circles.
+Panel (b): Resulting causal adjacency matrix $A^{k,j}$: $A^{k,j}=1$ (dark) for responses in the cluster with the larger centroid, and $A_{k,j}=0$ (white) otherwise.
+Example: the first row $A^{1,j}$ implies $x^{(1)}_{t+1}=f_{1}(x^{(1)}_{t},x^{(3)}_{t})$.
+Neural emulators.
+We learn two discrete physics-constrained neural models, with and without causal constraints, from the long trajectory $\mathbf{x}_{t}$.
+We denote these models as “Physics” and “Physics &\& Causal” emulators.
+Stationary statistics.
+We simulate a long trajectory of $T=10^{7}$ time steps with the neural models.
+We examine two stationary statistics: the invariant distribution (PDF) and the autocorrelation functions (ACFs).
+Both emulators faithfully reproduce these stationary statistics; see Sec.
+VII of the SM.
+Perturbed statistics: linear regime.
+We analyze the impulse response operator for the ensemble mean and variance.
+A small impulse perturbation $\delta x^{(j)}_{0}$ is applied at $t=0$, and the response of observables $A(x^{(k)}_{t})=x^{(k)}_{t}$ (mean) and $A(x^{(k)}_{t})=(x^{(k)}_{t}-\mu^{(k)}_{t})^{2}$ (variance) is measured, $\bm{\mu}_{t}$ representing the time-dependent mean (see SM Sec.
+VI).
+As a representative example, we perturb $x^{(6)}_{0}$ and report the self-response of $x^{(1)}_{t}$ in mean and variance (Fig.
+2(a,c)).
+Both emulators reproduce the response operator well, with the causally constrained model outperforming the non-causal one.
+This improvement is robust, as shown in SM Sec.
+VIII, where we report the time-dependent MSE of the responses $R^{k,j}_{t}$ aggregated over all pairs $j,k$.
+By contrast, we also trained a vanilla MLP network (no physics and no causal constraints).
+The vanilla model reproduced well the stationary distribution of the system, but suffered instability even under small perturbations (not shown).
+Figure 2: Perturbed statistics.
+Panel (a) and (c): linear response regime.
+Panel (a): Time-dependent, response in ensemble mean of variable $x^{(1)}_{t}$ to an impulse perturbation $\delta x^{(6)}_{0}$ imposed on $x^{(6)}_{0}$ at time $t=0$.
+Formally: $R^{1,6}_{t}=\delta\langle x^{(1)}_{t}\rangle/\delta x^{(6)}_{0}$.
+Panel (c): same as Panel (a) but for response in ensemble variance.
+Panel (b) and (d): nonlinear response regime.
+Panel (b): Time-dependent response of the ensemble mean of $x^{(3)}$ to a step function forcing.
+Forcing: step function $\mathbf{F}=(0,0,0,0,0,\sigma_{6})$ for $t\geq 0$, applied on the right-hand side of the numerical model and emulators.
+$\sigma_{6}$ is the standard deviation of $x^{(6)}$.
+Panel (d): same as Panel (b) but for the ensemble variance.
+Notably, an unconstrained vanilla MLP reproduces stationary statistics well but becomes unstable even under small perturbations (not shown).
+Perturbed statistics: nonlinear regime.
+We probe the nonlinear response by applying a large step forcing $\mathbf{F}$ to both the numerical model and the neural emulators.
+As in the linear analysis, the forcing acts on $x^{(6)}$, $\mathbf{F}=(0,0,0,0,0,\sigma_{6})$ for $t\geq 0$, where $\sigma_{6}$ is the standard deviation of $x^{(6)}$.
+We simulate an ensemble of $10^{5}$ realizations and measure the response of the ensemble mean and variance relative to the unperturbed stationary distribution.
+Results are shown in Figure 2(b,d).
+The numerical model exhibits a pronounced transient before relaxing to a new equilibrium after $\sim 40$ model time units, reflecting a shift of the attractor in state space.
+Both emulators remain stable under this strong forcing and predict the response well.
+The physics- and causally-constrained model generally outperformes the physics-only emulator.
+This improvement is robust as shown in Sec.
+IX where we report the time-dependent MSE of the responses aggregated over all perturbation experiments.
+IV Symmetry-broken Lorenz-96 system
+We define a symmetry-broken version of the Lorenz-96 (L96) [51] system as
+$\dot{x}_{j}=(x_{j+1}-x_{j-2})x_{j-1}-x_{j}+F_{j}+\sigma\xi_{j}(t),$ (6)
+with $j=1,\dots,J$ and periodic boundary conditions $x_{j}=x_{j+J}$.
+We set $J=20$.
+The time-indipendent forcing is defined as $F_{j}=F+\eta_{j}$, where $\eta_{j}\sim\mathcal{N}(0,\sigma_{F}^{2})$ with $\sigma_{F}=10$.
+Because the forcing includes a random component, we impose a minimum value of $4$ to maintain positivity.
+This site-dependent forcing (ranging from $4$ to $\sim 31$) breaks translational invariance and induces spatially heterogeneous turbulence.
+Physically, the localized forcing can be interpreted as topographic effects.
+The stochastic term $\sigma\xi_{j}(t)$, with $\sigma=5$, represents unresolved fast processes.
+The system in Eq.
+(6) serves as a prototype of a turbulent dynamical system, with strong mixing, exponential decay of autocorrelations, many unstable directions, and regular distributions even in the deterministic limit [58, 59].
+For reproducibility, the specific $F_{j}$ values are reported in SM Sec.
+X.
+We simulate a long trajectory of $T=10^{6}$ time steps with $dt=0.01$ ($10^{4}$ model time units).
+The 20-dimensional trajectory is treated as a discrete time series $\mathbf{x}_{t}=(x^{(1)}_{t},...,x^{(20)}_{t})$ from which we want to learn the discrete flow map in Eq.
 (1).
-Neural emulators.
-We learn two stochastic neural models from the long trajectory $\mathbf{x}(t)=[x_{1},...,x_{6}](t)$ of the CdV model: one with causal constraints and one without.
-The latter referred to as “vanilla” emulator.
-The causally constrained emulator leverages the Adjacency matrix $A_{k,j}$ inferred in Figure 1(c) to suppress spurious, non-causal terms through the new loss function in Eq.
+We infer the causal adjacency matrix following the protocol described in the previous Sections.
+As shown in Fig.
+3, the framework largely recovers the correct causal structure, including the periodic boundary conditions, despite the turbulent nonlinear dynamics.
+Two false negatives and one false positive are identified (red and blue in Fig.
+3).
+Because the inferred causal structure is imperfect, causal constraints are imposed through the robust loss in Eq.
 (4).
-The set $\mathcal{S}$ of pairs $(k,j)$ to be used in Eq.
-(10) is given by the zeros of $A_{k,j}$ in Figure 1(c) and equal to $\mathcal{S}=\{(1,2),(1,4),(1,5),(1,6),(2,5),(3,6),(4,1),(5,2),(6,3)\}$.
-Both neural emulators have been trained with the same hyper-parameters, the only difference being the addition of the regularization term as in Eq.
-(10) for the causally constrained model.
-Stationary statistics.
-Given the two emulators defined above, we simulate a long trajectory of $T=10^{7}$ time steps.
-We then examine two stationary statistics: the invariant probability distribution (PDF) and the autocorrelation functions (ACFs).
-The analysis is reported in Section VI of the SM: both the vanilla and the causally constrained emulator reproduce these stationary statistics well.
-Perturbed statistics: linear regime.
-We assess the ability of the emulators to reproduce perturbed statistics in the linear-response regime by analyzing the impulse response operator for the ensemble mean and variance, following [29].
-A small impulse perturbation $\delta x_{j}(0)$ is applied at $t=0$, and the response of observables $A(x_{k}(t))=x_{k}(t)$ (mean) and $A(x_{k}(t))=(x_{k}(t)-\mu_{k}(t))^{2}$ (variance) is measured, $\bm{\mu}(t)$ representing the time-dependent mean.
-Here, we perturb $x_{1}(0)$ and report the self-response of $x_{1}(t)$ in mean and variance (Fig.
-2).
-The vanilla emulator reproduces the response operator accurately, as expected given the large training dataset and Markovianity of the system [29].
-However, the causally constrained emulator yields a systematically improved representation of both mean and variance responses.
-This improvement persists for the full response operator $R_{k,j}(t)$ across all variables (see Figures 4–5 of the Supplemental Material).
-A quantitative summary of this improvement is provided by the total time-dependent MSE, aggregated over all $j$ and $k$, between the emulators and the ground truth, reported in Figure 6 of the SM.
-Figure 2: Perturbed statistics: linear regime.
-Self-response operator of mean and variance for the variable $x_{1}$.
-Panel (a): Time-dependent, response in ensemble mean of variable $x_{1}(t)$ to an impulse perturbation $\delta x_{1}(0)$ imposed on $x_{1}(0)$ at time $t=0$.
-Formally: $R_{1,1}(t)=\delta\langle x_{1}(t)\rangle/\delta x_{1}(0)$.
-Panel (b) same as Panel (a) but for response in ensemble variance.
-Responses are computed using an ensemble of $N_{e}=10^{6}$ members.
-The magnitude of the impulse perturbation is set to $10^{-1}\sigma_{1}$, $\sigma_{1}$ being the standard deviation of $x_{1}$.
-Labels: “Numerical” (numerical model), Model vanilla” (vanilla, unconstrained, emulator), and “Model causal” (causally constrained emulator).
-Perturbed statistics: nonlinear regime.
-We now probe the nonlinear response regime by applying a large step forcing $\mathbf{F}$ to both the numerical model and the neural emulators.
-As in the linear response analysis, the forcing is applied to $x_{1}$.
-The forced dynamics are given by $\dot{\mathbf{x}}=\mathbf{F}+\mathbf{f}(\mathbf{x})+\mathbf{\Sigma}\bm{\xi}(t)$, where $\mathbf{F}=(\sigma_{1},0,0,0,0,0)$ for $t\geq 0$, $\sigma_{1}$ being the std.
-dev.
-of the $x_{1}$ variable.
-The forcing amplitude is intentionally set very high to induce a strongly nonlinear response.
-We simulate an ensemble of $10^{5}$ forced realizations and measure the time-dependent response of the ensemble mean and variance relative to the unperturbed stationary distribution.
-The forcing alters the system’s attractor through time-dependent feedbacks across degrees of freedom.
-Figure 3 shows the resulting self-response of the marginal distribution of $x_{1}(t)$; responses of the full system are reported in Figures 7–8 of the Supplemental Material.
-The numerical model exhibits a pronounced transient response, relaxing toward a new equilibrium after approximately $40$ model time units.
-Therefore, the external forcing modifies the characteristics of the attractor in state space.
-The causally constrained emulator accurately reproduces these nonlinear forced responses, particularly in the ensemble mean.
-By contrast, the vanilla emulator becomes unstable after a short transient under the same forcing.
-This behavior is robust: when large step forcings are applied to each degree of freedom, the vanilla emulator becomes unstable in three out of six cases, whereas the causally constrained emulator remains stable for all forcings (see Section VIII of the SM).
-In general, despite the large forcing amplitude, we observe that the causally constrained emulator provides a good approximation of the forced responses in both ensemble mean and variance.
-Figure 3: Perturbed statistics: nonlinear regime.
-Time-dependent response of the ensemble mean and variance of $x_{1}$ to a strong step function forcing.
-Forcing: step function $\mathbf{F}=(\sigma_{1},0,0,0,0,0)$ for $t\geq 0$, applied on the right-hand side of the numerical model and emulators.
-$\sigma_{1}$ is the standard deviation of $x_{1}$.
-Responses are computed using an ensemble of $N_{e}=10^{5}$ members.
-Panel (a): response in ensemble mean.
-Panel (b): response in ensemble variance.
-Labels: “Numerical” (numerical model), “Model vanilla” (vanilla, unconstrained, emulator), and “Model causal” (causally constrained emulator).
-IV Discussion and conclusions
-We introduced a strategy that leverages the Fluctuation–Dissipation Theorem and score matching to (i) identify direct causal links from long trajectories of Markovian physical systems and (ii) impose this causal information as constraints in neural emulators.
-These constraints improve the ability of reduced-order neural emulators to capture responses to both weak and strong external perturbations, even if trained on unforced data alone.
-In practice, the choice of a Markovian, coarse-grained representation of an underlying high-dimensional dynamical system is itself a central challenge, as emphasized in classical work [79] and revisited in recent studies [72, 62, 29].
-However, given an appropriate choice of coarse-grained variables, the proposed constraints can be readily implemented in general neural network architectures.
-The framework is modular and can be combined with any data-driven estimator of the FDT operator, including recent proposals such as [61, 25], as well as with alternative causal discovery methods, for example [78].
-A natural next step is to extend this framework to observational climate data and to develop causally consistent, reduced-order neural models of the global atmosphere and ocean, including within emerging probabilistic architectures such as stochastic interpolants [18, 107, 77].
-This perspective also opens the possibility of combining causal constraints with complementary physical constraints, such as energy conservation [7], to jointly enforce stability and physical fidelity in neural emulators.
-Additionally, in realistic applications, data are often scarce and partially observed, leading to imperfect causal estimates.
-An important open direction is to assess how accurate such estimates must be - and whether approximate, quasi-Gaussian FDT operators may already suffice [68] - to construct imperfect emulators that nevertheless exhibit skill in responding to external perturbations.
-Finally, while motivated by climate dynamics, the proposed strategy applies to neural modeling of general complex turbulent dynamical systems in reduced coordinate spaces.
-In such settings, causal constraints provide a practical and readily implementable addition to general neural network architectures.
+Figure 3: Same as Figure 1, but for the L96 system defined in Eq.
+(6).
+Red squares indicate false negatives: true physical links erroneously labeled as noncausal.
+The blue square indicates a false positive: a spurious link erroneously labeled as causal.
+Example: the first row $A^{1,j}$ indicates a functional dependency $\mathbf{x}^{(1)}_{t+1}=f_{1}(x^{(19)}_{t},x^{(20)}_{t},x^{(1)}_{t},x^{(2)}_{t})$
+We then train a “Physics” and a “Physics &\& Causal” constrained neural emulators.
+Both reproduce the invariant density and autocorrelation functions (SM Sec.
+XA) and accurately capture the linear responses (not shown).
+Here we probe the nonlinear response regime by applying a Gaussian forcing centered at lattice site $j=10$,
+$G_{j}$ $=w\exp\!\left(-\frac{d(j,\mu)^{2}}{2\ell^{2}}\right),$ (7)
+$d(j,\mu)$ $=\min(|j-\mu|,\,J-|j-\mu|),$
+with $\mu=10$ and $\ell=3$.
+The distance function $d(j,\mu)$ is needed to account for the periodic boundary conditions.
+The forcing amplitude is determined by computing the total standard deviation $\sigma_{\text{Tot}}$ (which is on the order of the individual standard deviation $\sigma_{j}$ at each site $x^{(j)}$) and setting $w=10\sigma_{\text{Tot}}$.
+This constitutes a severe perturbation, an order of magnitude larger than the natural variability of any single degree of freedom.
+As depicted in Fig.
+4(top-left and bottom-left panels), this strong forcing significantly shifts the mean and variance of the attractor.
+The purely physics-constrained emulator already captures these statistical changes accurately.
+Augmenting the model with causal constraints yields a slightly improved representation of the response, particularly in the variance.
+Importantly, we emphasize that given the robust causal penalty in Eq.
+(4), the causal model maintains this high performance even if the imposed causal priors are imperfect.
+Finally, we also trained a vanilla MLP without physical or causal constraints: the unconstrained network was more difficult to train, requiring $\sim 10$ times more epochs to achieve a comparable loss.
+Although this baseline model reproduced the stationary distribution well and remained stable under the large forcing in Eq.
+(7) (not shown), it exhibited very large errors in the predicted response as also reported in Figure Fig.
+4.
+Figure 4: Response of the L96 system to a large Gaussian forcing [Eq.
+(7)] applied and held constant for $t\geq 0$.
+Top row: ensemble-mean response.
+Bottom row: ensemble-variance response.
+Left: ground truth.
+Second column: difference between the physics-constrained emulator and the truth.
+Third column: difference between the physics- and causally constrained emulator and the truth.
+Fourth column: difference between a completely unconstrained (no physics and no causal constraints) model and the truth.
+V Discussion and conclusions
+We have proposed a framework for physics and causally constrained reduced-order neural modeling of turbulent dynamical systems.
+The physics-constrained neural models are here formulated as discrete finite-time maps with strictly energy-conserving nonlinearities.
+This generalizes continuous-time conservation laws into a formulation inherently amenable to coarse-grained, discrete observational data.
+Building upon this architectural foundation, we demonstrated how causal information, derived via the fluctuation-dissipation theorem (FDT), can be practically enforced to suppress spurious functional dependencies within the neural model.
+Conservation laws alone do not uniquely determine the system’s dynamics; causal information provides additional constraints on admissible interactions.
+Crucially, while the energy-conserving architecture is prescribed a priori as a geometric prior, causal constraints must be inferred from data, requiring careful handling of estimation uncertainty.
+The resulting constrained models yield stable emulators that accurately capture stationary statistics and responses to both weak and strong external forcings, despite being trained exclusively on unforced data.
+A central challenge for practical data-driven applications remains the construction of a Markovian coarse-grained representation from high-dimensional spatiotemporal data, as emphasized in classical theory [70] and explored in recent work [63, 55, 23].
+However, given an approximate Markovian, coarse-grained representation of a spatiotemporal system, informed by physics and domain knowledge, the proposed constrained discrete neural models can be readily adopted.
+A natural extension of this work is its application to observational climate data to construct reduced-order emulators of the global atmosphere and ocean.
+More broadly, while our test cases are motivated by climate dynamics, this methodology provides a practical strategy for neural modeling of general complex turbulent systems in reduced coordinate spaces.
 Code availability
@@
 F.F.
-is grateful to Rory Basinski and Matthieu Blanke for many enriching discussions on this subject.
+is grateful to Rory Basinski, Matthieu Blanke and Andre Souza for enriching discussions on this subject.
 Supplemental Material for:
-Causally constrained reduced-order neural climate emulators via response theory
+Physics and causally constrained discrete-time neural models
+of turbulent dynamical systems
 Fabrizio Falasca and Laure Zanna
@@
 Fabrizio Falasca Laure Zanna
-Figure 1: Schematic of the proposed framework.
-Step 1: From a long trajectory of a Markovian system, the Fluctuation–Dissipation Theorem is used to infer the system’s causal structure.
-Step 2: This inferred structure is then imposed as a soft constraint during neural-network training, yielding causally constrained emulators.
-The example shows time series from the Charney–DeVore model (Eq.
-(3) in the main text) and the discovered causal graph, but the approach applies to general stochastic dynamical systems.
-I Fluctuation-Dissipation Theorem (FDT) from data: quasi-Gaussian and score matching
+I Physics and causally constrained models
+I.1 Previous work on continuous models
+The models considered in this work build on the following abstract representation of turbulent dynamical systems:
+$\dot{\mathbf{x}}=\mathbf{F}+\mathbf{A}\mathbf{x}+\mathbf{B}(\mathbf{x},\mathbf{x})+\mathbf{\Sigma}\bm{\xi}(t).$ (1)
+Such a functional form, characterized by energy-conserving quadratic nonlinearities $\mathbf{B}(\mathbf{x},\mathbf{x})$, typically arises when turbulent dynamical systems are projected onto a reduced orthogonal basis [50, 56].
+Here $\mathbf{F}$ represents an external forcing, $\mathbf{A}$ a linear operator, and $\mathbf{\Sigma}\bm{\xi}(t)$ a stochastic forcing parametrizing unresolved degrees of freedom.
+Majda and collaborators [64, 65] first emphasized the importance of constraining nonlinear regression models inferred from time series so that the nonlinear terms are energy conserving.
+A practical data-driven implementation of such constraints was later proposed by Kondrashov et al.
+[44].
+While most work has focused on constraining the nonlinear component $\mathbf{B}(\mathbf{x},\mathbf{x})$, the linear operator can also be physically constrained.
+Indeed, many physical systems of the form in Eq.
+(1) admit a decomposition of the linear operator $\mathbf{A}$ into dispersive and dissipative components, $\mathbf{A}=\mathbf{L}+\mathbf{D}$, where $\mathbf{L}$ is skew-symmetric ($\mathbf{L}^{\mathrm{T}}=-\mathbf{L}$, representing dispersion) and $\mathbf{D}$ is symmetric negative-definite ($\mathbf{D}^{\mathrm{T}}=\mathbf{D}<0$), representing dissipation [59, 58].
+Recent work by Sanderse and collaborators [84, 92] has further highlighted a more general representation of nonlinear interactions using skew-symmetric operators of the form $\mathbf{S}(\mathbf{x})\mathbf{x}$.
+This motivates a more general formulation:
+$\dot{\mathbf{x}}=\mathbf{F}+\mathbf{A}\mathbf{x}+\mathbf{S}(\mathbf{x})\mathbf{x}+\mathbf{\Sigma}(\mathbf{x})\bm{\xi}(t),$ (2)
+where the linear operators $\mathbf{L}$ and $\mathbf{D}$ satisfy the constraints described above, and the nonlinear operator is required to be skew-symmetric, $\mathbf{S}(\mathbf{x})^{\mathrm{T}}=-\mathbf{S}(\mathbf{x})$.
+Additionally, we impose $\mathbf{S}(0)=0$, ensuring that $\mathbf{S}(\mathbf{x})$ contains only nonlinear contributions.
+Energy conservation of the skew-symmetric nonlinear operator $\mathbf{S}(\mathbf{x})$.
+For completeness, we show how the skew-symmetric nonlinear operator leads to energy conservation. We define the system “energy” as the quadratic functional
+$E=(1/2)\mathbf{x}^{\mathrm{T}}\mathbf{x},$
+whose time derivative is
+$\frac{dE}{dt}=\frac{d}{dt}(\frac{1}{2}\mathbf{x}^{\mathrm{T}}\mathbf{x})=\mathbf{x}^{\mathrm{T}}\frac{d\mathbf{x}}{dt}.$
+Isolating the nonlinear contribution in Eq. (2),
+$\frac{dE}{dt}=\mathbf{x}^{\mathrm{T}}\mathbf{S}(\mathbf{x})\mathbf{x}.$
+Since $\mathbf{S}(\mathbf{x})$ is skew-symmetric, $\mathbf{S}(\mathbf{x})^{\mathrm{T}}=-\mathbf{S}(\mathbf{x})$, it follows that
+$\mathbf{x}^{\mathrm{T}}\mathbf{S}(\mathbf{x})\mathbf{x}=0,$
+leading to energy conservation $dE/dt=0$ of the nonlinear terms. We emphasize that this conservation property applies only to the nonlinear term.
+I.2 Proposed discrete-time models
+The structural constraints discussed above are formally derived for continuous-time dynamical systems in the limit $dt\to 0$.
+However, neural emulators are often inferred from discrete time series.
+In practice, given $T$ observations $\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{T}$ sampled at a fixed interval, this effective time step reflects (non-unique) choices of temporal coarse-graining rather than a mere numerical discretization parameter [23].
+It is therefore desirable to design a model trained as a discrete-time predictor $\mathbf{x}_{t+1}\approx\mathbf{f}(\mathbf{x}_{t})$, while still preserving the fundamental symmetries showcased above.
+In particular, we focus on the crucial constraint of energy conservation within the nonlinear term.
+In the discrete case, we require that the nonlinear update conserves the $L_{2}$ norm of the state it acts on; i.e., the “energy” of the system at time $t$ is defined as $\|\mathbf{x}_{t}\|^{2}$.
+As introduced in the main text, we propose the following functional form:
+$\mathbf{v}_{t}$ $=\mathbf{Q}(\mathbf{x}_{t})\mathbf{x}_{t},$ (3)
+$\mathbf{x}_{t+1}$ $=\mathbf{F}+\mathbf{M}\mathbf{v}_{t}+\mathbf{\Sigma}\bm{\xi}_{t},$
+where $\mathbf{F}$ and $\mathbf{\Sigma}\bm{\xi}_{t}$ represent the effective deterministic and stochastic forcings, arising from integrating over the discrete finite time step (not to be confused with their continuous counterparts in Eq.
+(2)).
+The nonlinear operator $\mathbf{Q}(\mathbf{x}_{t})$ is constrained to be strictly orthogonal ($\mathbf{Q}^{\top}\mathbf{Q}=\mathbf{I}$).
+The isolated nonlinear term $v_{t}=\mathbf{Q}(\mathbf{x}_{t})\mathbf{x}_{t}$ contributes no net growth or decay to the energy budget, strictly preserving the $L_{2}$-norm of the state:
+$\|\mathbf{v}_{t}\|^{2}=\|\mathbf{Q}(\mathbf{x}_{t})\mathbf{x}_{t}\|^{2}=\mathbf{x}_{t}^{\top}\mathbf{Q}^{\top}(\mathbf{x}_{t})\mathbf{Q}(\mathbf{x}_{t})\mathbf{x}_{t}=\mathbf{x}_{t}^{\top}\mathbf{I}\mathbf{x}_{t}=\|\mathbf{x}_{t}\|^{2}.$
+Geometrically, the full deterministic update rule acts as two sequential operations: a rotation in state space, $\mathbf{v}_{t}=\mathbf{Q}(\mathbf{x}_{t})\mathbf{x}_{t}$, followed by a linear transformation $\mathbf{M}\mathbf{v}_{t}$.
+The overall energy growth or decay of the system is exclusively controlled by the singular values of the linear operator $\mathbf{M}$.
+Additional constraints can, in principle, be applied to guarantee global dissipative properties by strictly bounding the maximum singular value of $\mathbf{M}$.
+We view these strict constraints on the linear operator $\mathbf{M}$ as optional, and potentially reducing the flexibility of the resulting data-driven model.
+We therefore leave potential constraints on $\mathbf{M}$ to future work.
+I.2.1 Details on the emulator implementation
+Given a $n$-dimensional discrete dynamical system encoded in as a long trajectory $\mathbf{x}\in\mathbb{R}^{n,T}$, with length $T$, we aim in fitting as:
+[…]
+where the deterministic drift is
+$\mathbf{f}(\mathbf{x}_{t})=\mathbf{F}+\mathbf{M}\mathbf{S}(\mathbf{x}_{t})\mathbf{x}_{t}.$
+Fitting the deterministic drift $\mathbf{f}(\mathbf{x}_{t})$.
+We fit the deterministic dynamics such that $\mathbf{x}_{t+1}\approx\mathbf{f}(\mathbf{x}_{t})$. We do so by minimizing the mean squared error (MSE) over the trajectory:
+$\mathcal{L}=\text{MSE}(\mathbf{x}_{t+1},\mathbf{f}(\mathbf{x}_{t})).$ (5)
+  • All terms comprising the discrete map, $\mathbf{f}(\mathbf{x}_{t})=\mathbf{F}+\mathbf{M}\mathbf{Q}(\mathbf{x}_{t})\mathbf{x}_{t}$, are trained jointly via gradient descent.
+The primary challenge in this joint optimization is the initialization step.
+To resolve this, the linear components $\mathbf{F}$ and $\mathbf{M}$ are first pre-computed using Ordinary Least Squares (OLS) to capture the baseline linear dynamics.
+This deterministic “first guess” serves as the initialization for the network parameters.
+  • The nonlinear operator $\mathbf{Q}(\cdot)$ is parameterized by a Multi-Layer Perceptron (MLP) with a SiLU activation function.
+To strictly enforce orthogonality ($\mathbf{Q}^{\mathrm{T}}\mathbf{Q}=\mathbf{I}$) across all forward passes, we employ the following architectural construction:
+    • The MLP maps a state $\mathbf{x}_{t}\in\mathbb{R}^{n}$ to n(n1)/2n(n-1)/2 independent scalar values.
+    • These values populate the upper-triangular elements of a matrix, which is then used to construct an exact skew-symmetric matrix $\mathbf{S}(\mathbf{x}_{t})=-\mathbf{S}^{\mathrm{T}}(\mathbf{x}_{t})$.
+    • We map the skew-symmetric $\mathbf{S}(\mathbf{x}_{t})$ to an orthogonal operator via the matrix exponential, $\mathbf{Q}(\mathbf{x}_{t})=\exp(\mathbf{S}(\mathbf{x}_{t}))$, or the computationally efficient Cayley transform, $\mathbf{Q}(\mathbf{x}_{t})=(\mathbf{I}-\mathbf{S}(\mathbf{x}_{t}))^{-1}(\mathbf{I}+\mathbf{S}(\mathbf{x}_{t}))$.
+In the main text we use matrix exponential as default option.
+    Crucially, the weights of the final linear layer of the MLP are initialized to exactly zero.
+Consequently, at the onset of training, the network outputs $\mathbf{S}(\mathbf{x}_{t})=\mathbf{0}$, which yields $\mathbf{Q}(\mathbf{x}_{t})=\mathbf{I}$.
+This initialization strategy guarantees that the joint training step begins precisely at the stable OLS baseline before gradually introducing the parameterized nonlinear dynamics.
+Data standardization.
+Data standardization requires careful treatment, as the strict physical constraint of energy conservation (the $L_{2}$ norm) depends on the geometry of the original variables and is not generally preserved in a shifted and scaled coordinate space.
+To ensure numerical stability during optimization, the neural network emulator is trained on standardized variables, obtained by transforming each time series in $\mathbf{x}_{t}$ to zero mean and unit variance:
+$\tilde{\mathbf{x}}_{t}=\mathbf{\Sigma}_{\text{std}}^{-1}(\mathbf{x}_{t}-\bm{\mu}),$
+where $\bm{\mu}$ contains the empirical means and $\mathbf{\Sigma}_{\text{std}}$ is a diagonal matrix of standard deviations.
+The orthogonal constraint on the nonlinear operator $\mathbf{Q}$ is formulated to conserve energy exclusively in the physical space.
+Accordingly, while the neural network generating $\mathbf{Q}(\tilde{\mathbf{x}}_{t})$ takes standardized inputs, the resulting rotation acts on the unstandardized physical state to yield an intermediate rotated state $\mathbf{v}_{t}$:
+$\mathbf{v}_{t}=\mathbf{Q}(\tilde{\mathbf{x}}_{t})\big(\mathbf{\Sigma}_{\text{std}}\tilde{\mathbf{x}}_{t}+\bm{\mu}\big).$
+We note that $\mathbf{Q}(\tilde{\mathbf{x}}_{t})$ still acts on the standardized states $\tilde{\mathbf{x}}_{t}$ as it usually facilitate training.
+To apply the linear dynamics, this physically rotated state must be mapped back into the standardized coordinate space, effectively removing the physical offsets:
+$\tilde{\mathbf{v}}_{t}=\mathbf{\Sigma}_{\text{std}}^{-1}(\mathbf{v}_{t}-\bm{\mu}).$
+Finally, the linear operator $\mathbf{M}$ and the effective deterministic forcing $\mathbf{F}$ (both parameterized in the standardized space) act on this standardized rotated state to yield the full discrete update:
+$\tilde{\mathbf{x}}_{t+1}=\mathbf{F}+\mathbf{M}\tilde{\mathbf{v}}_{t}.$
+This careful separation guarantees that the nonlinear interactions remain strictly energy-conserving with respect to the original physical variables, while the linear dissipation (and the neural network optimization) operate in a well-conditioned standardized space.
+Finally, we note that by virtue of standardizing the state variables to zero mean, the constant forcing vector $\mathbf{F}$ is essentially zero at initialization and remains negligibly small during training.
+Fitting the noise term.
+Once the deterministic discrete map $\mathbf{f}$ is estimated, we close the model with a stochastic forcing.
+We compute the residuals of the deterministic dynamics as $\mathbf{r}_{t}=\mathbf{x}_{t+1}-\mathbf{f}(\mathbf{x}_{t})$, with $\mathbf{r}_{t}\in\mathbb{R}^{n}$, and construct a stochastic parameterization based on these residuals.
+As the primary focus of this work is on constraining the deterministic drift, we adopt the simplest choice of additive diagonal noise: the diagonal matrix $\mathbf{\Sigma}$ with each entry computed as the standard deviation of the residual component $r^{(i)}_{t}$.
+More general stochastic parameterizations, including state-dependent and correlated noise, can be readily constructed from $\mathbf{r}_{t}$; we refer to Appendix C of [23] for details.
+Finally, the proposed physics-constrained architectures can be naturally combined with multilevel regression strategies [44] to account for non-Markovian memory effects arising from unresolved scales, which we leave for future work.
+I.3 Causal constraints
+The physics-constrained networks detailed above can be augmented via a causal regularization term by adding a quadratic penalty to the MSE loss:
+$\mathcal{L}$ $=\mathcal{L}_{\text{MSE}}+\mathcal{L}_{\text{Causal}}$ (6)
+$=\text{MSE}(\mathbf{x}_{t+1},\mathbf{f}(\mathbf{x}_{t}))+\lambda\sum_{(k,j)\in\mathcal{S}}\left(\frac{\partial f_{k}}{\partial x^{(j)}}\right)^{2},$
+where the gradients $\frac{\partial f_{k}}{\partial x^{(j)}}$ are efficiently computed during training via automatic differentiation [71].
+In the ideal case of recovering the exact causal graph, the purely quadratic penalty in Eq.
+(6) suffices to enforce the correct sparsity pattern in the discrete map’s Jacobian.
+However, in high-dimensional systems with finite data, causal estimates are inherently uncertain, and False Negatives (i.e.
+true causal couplings $x^{(j)}\rightarrow x^{(k)}$ mistakenly classified as non-causal) may occur.
+Under a strict quadratic penalty, such errors would severely bias the learned dynamics by artificially suppressing the required gradients.
+To mitigate this, we propose a robust, capped penalty in the main text:
+$\mathcal{L}_{\text{Causal}}^{\text{Robust}}=\lambda\sum_{(k,j)\in\mathcal{S}}\min\!\left(\left(\frac{\partial f_{k}}{\partial x^{(j)}}\right)^{2},\gamma\right).$ (7)
+This capped formulation strictly bounds the penalty.
+Once the squared gradient reaches the threshold $\gamma$, the derivative of the regularizer with respect to the network weights vanishes.
+Consequently, if the MSE loss strongly dictates that a specific coupling is necessary to predict $\mathbf{x}_{t+1}$ (i.e., a mistakenly excluded interaction), the MSE term will easily overcome the bounded penalty.
+False Positives, by contrast, do not induce structural bias, as they are simply excluded from the set $\mathcal{S}$ and thus unpenalized.
+I.3.1 Heuristic for the “cap” parameter $\gamma$
+We propose a heuristic to estimate the cap parameter $\gamma$ directly from the discrete impulse response operator $R^{k,j}_{1}$ obtained via the FDT.
+The impulse response $R^{k,j}_{t}$ measures the change in variable $x^{(k)}$ at time step $t$ following a small perturbation applied to $x^{(j)}$ at step 0.
+As described in the main text, direct causal links $x^{(j)}\rightarrow x^{(k)}$ are identified from the unit-step responses $R^{k,j}_{1}$ by: (i) applying a logarithmic transformation $\ln|R^{k,j}_{1}|$ for $k\neq j$, (ii) partitioning these values using $k$-means clustering ($k=2$), and (iii) assigning a causal link to the cluster associated with the larger centroid.
+Let $\tilde{R}^{k,j}_{1}$ denote the subset of responses belonging to this cluster, corresponding to statistically significant (non-spurious) interactions.
+We define the cap parameter $\gamma$ as a low quantile of these squared significant responses; in our experiments, we use the $0.1$ quantile.
+This choice establishes a strict lower bound on the magnitude of reliably detectable causal interactions.
+If during training the MSE term drives a penalized gradient beyond this threshold, if the MSE term forces a penalized gradient to exceed this threshold, i.e.
+$\left(\partial f_{k}/\partial x^{(j)}\right)^{2}\geq\gamma$), this indicates that the interaction is likely inconsistent with the assumed absence of a causal link (i.e.
+we are dealing with a False Negative).
+In this regime, the capped loss effectively deactivates the penalty, allowing the dynamics to be governed by the MSE term rather than by a potentially incorrect causal constraint.
+I.3.2 Implementation details for the causal constraint
+Enforcing the causal penalty defined in Eq.
+(7) requires computing the Jacobian of the deterministic forward map in Eq.
+(3).
+However, backpropagation through $\mathbf{Q}(\mathbf{x}_{t})=\exp(\mathbf{S}(\mathbf{x}_{t}))$ is computationally expensive.
+Therefore to enforce the causal constraints efficiently, we construct a first-order proxy for the forward map through the skew-symmetric infinitesimal generator $\mathbf{S}(\mathbf{x}_{t})$.
+By expanding the orthogonal map to first order, $\mathbf{Q}(\mathbf{x}_{t})\approx\mathbf{I}+\mathbf{S}(\mathbf{x}_{t})$, the deterministic update $\mathbf{M}\mathbf{Q}(\mathbf{x}_{t})\mathbf{x}_{t}$ can be evaluated for the purpose of the Jacobian penalty as:
+$\tilde{\mathbf{f}}(\mathbf{x}_{t})=\mathbf{M}\mathbf{x}_{t}+\mathbf{M}\mathbf{S}(\mathbf{x}_{t})\mathbf{x}_{t}.$ (8)
+Constant forcing terms are ignored in this proxy, as their derivative with respect to the state is zero.
+Crucially this proxy is evaluated strictly to compute the gradients $\partial\tilde{f}_{k}/\partial x^{(j)}$ for the regularization term $\mathcal{L}_{\text{Causal}}^{\text{Robust}}$.
+The primary state prediction $\mathbf{x}_{t+1}$, and its associated MSE loss, are always computed using the exact, energy-conserving orthogonal map $\mathbf{M}\mathbf{Q}(\mathbf{x}_{t})\mathbf{x}_{t}$.
+This dual-path approach allows us to bypass the differentiation of the matrix exponential entirely.
+II Mathematical origin of the discrete model and a few examples
+The data-driven model proposed in this work is formulated and solved as discrete-time map and it is fundamentally designed to simulate coarse-grained, effective dynamics directly from data.
+To establish the theoretical grounding of this discrete formulation, we first demonstrate its mathematical connection to continuous-time systems.
+We further reformulate two illustrative dynamical systems via the proposed formulation.
+Finally, we test the framework by learning the effective dynamics of severely sub-sampled trajectories of the Charney-DeVore dynamical systems considered in the main text, targeting a regime where data-driven continuous models can fail due to numerical instability.
+In summary, this Section targets the following objectives:
+  • Formulation in the continuous limit.
+To build mathematical intuition, we first examine the limit of small time steps ($\Delta t\to 0$).
+We show that the general discrete architecture proposed in the main text arises naturally from the continuous equations via a Lie–Trotter splitting procedure [7].
+We then further highlight this connection with practical examples by focusing on two canonical systems:
+    • Deterministic “warm-up” with Lorenz-63.
+We reformulate the Lorenz-63 equations [50] into a structure consistent with our discrete emulator, yielding an elegant geometric interpretation of its chaotic dynamics.
+    • Stochastic triad model. We extend this geometric intuition to stochastic processes by reformulating a known triad model with energy-conserving nonlinearities into the proposed framework.
+  • Effective coarse-grained dynamics (Charney–DeVore).
+Finally, we demonstrate the primary use case of our method: stable learning from severely coarse-grained data.
+Using the Charney–DeVore model presented in the main text, we train the framework on trajectories subsampled every 100 and 500 integration steps.
+In this limit, both standard numerical integrators and physics-constrained, continuous data-driven models inevitably go unstable.
+By contrast, our learned discrete map acts as a stable emulator of the effective finite-time dynamics.
+II.1 Connection to the continuous case: general stochastic processes
+We consider the general abstract formulation arising when many turbulent flows are projected onto orthogonal bases, as discussed in the main text and at the start of this SM.
+We write this formulation as:
+$d\mathbf{x}=[\mathbf{F}+\mathbf{A}\mathbf{x}+\mathbf{S}(\mathbf{x})\mathbf{x}]dt+\mathbf{\Sigma}d\mathbf{W},$ (9)
+where $\mathbf{W}$ is a standard Wiener process and $\mathbf{S}(\mathbf{x})$ is strictly skew-symmetric, leading to energy conservation of the nonlinear term.
+For details on the various terms, see Section I.1.
+To map this continuous system to our proposed discrete-time architecture over a time step $\Delta t$, we apply a first-order operator splitting [7].
+This splitting procedure consists of two steps: we first decompose the dynamics into a purely nonlinear, energy-conserving step, followed by a linear, forced, and stochastic step.
+Step (a): Nonlinear rotation.
+We first isolate the nonlinear advection term:
+$d\mathbf{x}^{(1)}=\mathbf{S}(\mathbf{x}^{(1)})\mathbf{x}^{(1)}dt.$ (10)
+We approximate the solution over the interval $t,t+\Delta t]$ by freezing the state dependence of $\mathbf{S}$ at the beginning of the interval.
+Using the initial condition $\mathbf{x}^{(1)}(t)=\mathbf{x}_{t}$, this yields the approximate nonlinear update:
+$\mathbf{v}_{t}=\exp\big(\Delta t\mathbf{S}(\mathbf{x}_{t})\big)\mathbf{x}_{t}\equiv\mathbf{Q}(\mathbf{x}_{t})\mathbf{x}_{t}.$ (11)
+Because $\mathbf{S}(\mathbf{x}_{t})$ is skew-symmetric, its matrix exponential $\mathbf{Q}(\mathbf{x}_{t})$ is strictly orthogonal ($\mathbf{Q}^{\mathrm{T}}\mathbf{Q}=\mathbf{I}$).
+Thus, this step represents a pure rotation in state space that strictly preserves the $L_{2}$-norm (energy).
+Analytically, this step incurs a $\mathcal{O}(\Delta t^{2})$ truncation error.
+Step (b): Linear, forced, and stochastic flow.
+Next, we use the rotated intermediate state $\mathbf{v}_{t}$ as the initial condition for the remaining terms over the same time interval, leading to the linear SDE:
+$d\mathbf{x}^{(2)}=\big[\mathbf{A}\mathbf{x}^{(2)}+\mathbf{F}\big]dt+\mathbf{\Sigma}d\mathbf{W}_{t},\quad\text{with }\mathbf{x}^{(2)}(t)=\mathbf{v}_{t}.$ (12)
+Its exact solution over one time step is:
+$\mathbf{x}_{t+1}=e^{\mathbf{A}\Delta t}\mathbf{v}_{t}+\int_{0}^{\Delta t}e^{\mathbf{A}(\Delta t-s)}\mathbf{F}\,ds+\int_{0}^{\Delta t}e^{\mathbf{A}(\Delta t-s)}\mathbf{\Sigma}\,d\mathbf{W}_{s}.$ (13)
+We identify the following terms:
+  • Linear Operator: $\mathbf{M}=e^{\mathbf{A}\Delta t}$.
+The term $e^{\mathbf{A}\Delta t}\mathbf{v}_{t}$ in Eq.
+(12) then represents the sequential update $\mathbf{M}\mathbf{Q}(\mathbf{x_{t}})\mathbf{x_{t}}$ considered in the previous Sections.
+  • Effective Deterministic Forcing: $\mathbf{F}_{\text{discrete}}=\int_{0}^{\Delta t}e^{\mathbf{A}(\Delta t-s)}\mathbf{F}\,ds$.
+Assuming $\mathbf{A}$ is invertible and that $\mathbf{F}$ is a constant forcing, the solution of this integral is:
+    $\mathbf{F}_{\text{discrete}}=\mathbf{A}^{-1}\big(e^{\mathbf{A}\Delta t}-\mathbf{I}\big)\mathbf{F}=\mathbf{A}^{-1}(\mathbf{M}-\mathbf{I})\mathbf{F}.$ (14)
+  • Effective Stochastic Forcing: $\mathbf{\Sigma}_{\text{discrete}}\bm{\xi}_{t}=\int_{0}^{\Delta t}e^{\mathbf{A}(\Delta t-s)}\mathbf{\Sigma}\,d\mathbf{W}_{s}$.
+The covariance matrix of this process can be written as:
+    $\mathbf{C}=\mathbf{\Sigma}_{\text{discrete}}\mathbf{\Sigma}_{\text{discrete}}^{\mathrm{T}}=\int_{0}^{\Delta t}e^{\mathbf{A}(\Delta t-s)}\mathbf{\Sigma}\mathbf{\Sigma}^{\mathrm{T}}e^{\mathbf{A}^{\mathrm{T}}(\Delta t-s)}\,ds,$
+    where $\mathbf{\Sigma}_{\text{discrete}}$ is the lower-triangular Cholesky factor of $\mathbf{C}$, and $\bm{\xi}_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$.
+Importantly, $\mathbf{C}$ satisfies the Lyapunov matrix equation [91]:
+    $\mathbf{A}\mathbf{C}+\mathbf{C}\mathbf{A}^{\mathrm{T}}=\mathbf{M}\mathbf{\Sigma}\mathbf{\Sigma}^{\mathrm{T}}\mathbf{M}^{\mathrm{T}}-\mathbf{\Sigma}\mathbf{\Sigma}^{\mathrm{T}},\qquad\mathbf{M}=e^{\mathbf{A}\Delta t}.$ (15)
+    Thus, $\mathbf{C}$ can be identified by numerically solving the Lyapunov equation above. The term $\mathbf{\Sigma}_{\text{discrete}}$ is then extracted via Cholesky decomposition of $\mathbf{C}$.
+Combining these steps yields a split numerical approximation of the forward dynamics:
+$\mathbf{x}_{t+1}=\mathbf{F}_{\text{discrete}}+\mathbf{M}\mathbf{Q}(\mathbf{x}_{t})\mathbf{x}_{t}+\mathbf{\Sigma}_{\text{discrete}}\bm{\xi}_{t}.$ (16)
+Which correspond to the same formulation in Eq.
+(1) in the main text.
+This derivation is formally first-order accurate, with a local $\mathcal{O}(\Delta t^{2})$ splitting error and it precisely mirrors the discrete model employed in the main text.
+In contrast, the data-driven formulation does not impose any explicit requirement on the temporal resolution and is tailored to fit the effective dynamics of the system at coarse-grained time scales, which can differ substantially from the underlying continuous dynamics.
+At the same time, it enforces the resolution-independent orthogonality constraint of the nonlinear update $\mathbf{Q}(\mathbf{x}_{t})\mathbf{x}_{t}$, guaranteeing that the nonlinear term cannot induce spurious energy growth or decay.
+II.1.1 Deterministic regime. Pedagogical example with the Lorenz-63 model
+We now show how to reformulate the Lorenz-63 system [50] as in (16).
+In this simple first case, the dynamics will be governed by the update $\mathbf{x}_{t+1}=\mathbf{M}\mathbf{Q}(\mathbf{x}_{t})\mathbf{x}_{t}$.
+The continuous equations governing the state $\mathbf{x}=(x,y,z)$ are given by:
+$\dot{x}$ $=\sigma(y-x),$ (17)
+$\dot{y}$ $=x(\rho-z)-y,$
+$\dot{z}$ $=xy-\beta z.$
+We can exactly decompose these dynamics into a constant linear operator $\mathbf{A}$ and a state-dependent, strictly skew-symmetric matrix $\mathbf{S}(\mathbf{x})$:
+$\dot{\mathbf{x}}=\underbrace{\begin{bmatrix}-\sigma&\sigma&0\\ \rho&-1&0\\ 0&0&-\beta\end{bmatrix}}_{\mathbf{A}}\begin{bmatrix}x\\ y\\ z\end{bmatrix}+\underbrace{\begin{bmatrix}0&0&0\\ 0&0&-x\\ 0&x&0\end{bmatrix}}_{\mathbf{S}(\mathbf{x})}\begin{bmatrix}x\\ y\\ z\end{bmatrix}$.
+(18)
+Here, the constant linear matrix $\mathbf{A}$ encapsulates the system’s local instabilities as well as dissipation.
+The nonlinear interactions are entirely isolated within $\mathbf{S}(\mathbf{x})$.
+Because $\mathbf{S}(\mathbf{x})^{\top}=-\mathbf{S}(\mathbf{x})$, the nonlinear term is strictly energy-preserving, satisfying $\mathbf{x}^{\top}\mathbf{S}(\mathbf{x})\mathbf{x}=0$.
+The application of Lie–Trotter operator splitting [7] as explained in the above Section II.1, leads to
+$\mathbf{x}_{t+1}=\mathbf{M}\mathbf{Q}(\mathbf{x}_{t})\mathbf{x}_{t},$ (19)
+where $\mathbf{x}_{t}=(x_{t},y_{t},z_{t})$, $\mathbf{M}=\exp(\Delta t\mathbf{A})$, and $\mathbf{Q}(\mathbf{x}_{t})=\exp(\Delta t\mathbf{S}(\mathbf{x}_{t}))$.
+Given that the infinitesimal generator $\mathbf{S}(\mathbf{x}_{t})$ is skew-symmetric, its matrix exponential $\mathbf{Q}(\mathbf{x}_{t})$ is strictly orthogonal.
+The nonlinear update matrix $\mathbf{Q}(\mathbf{x}_{t})$ of the Lorenz-63 system can be calculated analytically and it is equal to:
+$\mathbf{Q}(\mathbf{x}_{t})=\exp(\Delta t\mathbf{S}(\mathbf{x}_{t}))=\begin{bmatrix}1&0&0\\ 0&\cos(x_{t}\Delta t)&-\sin(x_{t}\Delta t)\\ 0&\sin(x_{t}\Delta t)&\cos(x_{t}\Delta t)\end{bmatrix}$ (20)
+The linear operator $\mathbf{M}$ can be computed numerically once a $\Delta t$ is specified.
+For a chosen time step (e.g., $\Delta t=0.01$) and standard Lorenz parameters ($\sigma=10,\rho=28,\beta=8/3$), this evaluates to a constant matrix:
+$\mathbf{M}=\exp(\Delta t\mathbf{A})\underset{\Delta t=0.01}{=}\begin{bmatrix}0.917925&0.0951229&0\\ 0.266344&1.00354&0\\ 0&0&0.973686\end{bmatrix}$ (21)
+Given an initial condition $\mathbf{x}_{t=0}$, the discrete mapping explicitly factors the dynamics into two sequential geometric operations.
+First, the state-dependent orthogonal update $\mathbf{v}_{t}=\mathbf{Q}(\mathbf{x}_{t})\mathbf{x}_{t}$ applies an energy-preserving rotation strictly in the $y$-$z$ plane.
+The rotation angle $\theta=x_{t}\Delta t$ is state-dependent and dictated by the instantaneous intensity of the $x_{t}$ coordinate, so nearby trajectories experience different rotations.
+Second, the linear transformation $\mathbf{x}_{t+1}=\mathbf{M}\mathbf{v}_{t}$ stretches and contracts the rotated vector $\mathbf{v}_{t}$.
+Crucially, unlike a purely dissipative operator, in the Lorenz system, $\mathbf{M}$ is not contractive in its norm: it expands some directions while contracting others, making it a special case in respect to the class of systems considered in this study.
+A long integration of this map reproduces the Lorenz attractor with high-fidelity, see Figure 1. The code to reproduce such results can be found in the code availability Section in the main text.
+Figure 1: Lorenz attractor simulated with the sequential mapping scheme $x_{t+1}=\mathbf{M}\mathbf{Q}(\mathbf{x}_{t})\mathbf{x}_{t}$.
+II.1.2 Stochastic regime. Pedagogical example with triad models
+We now consider the stochastic triad model proposed in [58] and formulated as:
+$dx_{1}$ $=L_{2}x_{3}-L_{3}x_{2}-d_{1}x_{1}+B_{1}x_{2}x_{3}+F_{1}+\sigma_{1}dW_{1},$ (22)
+$dx_{2}$ $=L_{3}x_{1}-L_{1}x_{3}-d_{2}x_{2}+B_{2}x_{1}x_{3}+F_{2}+\sigma_{2}dW_{2}$
+$dx_{3}$ $=L_{1}x_{2}-L_{2}x_{1}-d_{3}x_{3}+B_{3}x_{1}x_{2}+F_{3}+\sigma_{3}dW_{3}.$
+This model serves as a fundamental building block for complex turbulent dynamical systems, as three-dimensional Galerkin truncations of various fluid equations naturally yield this structural form (see Chapter 2 of Ref.
+[59]).
+The linear operator $\mathbf{A}$ can be decomposed in $\mathbf{A}=\mathbf{L}+\mathbf{D}$, where $\mathbf{L}$ is a skew-symmetric matrix representing dispersion, and $\mathbf{D}$ is a symmetric negative-definite matrix representing dissipation.
+In the specific case in Eq.
+(22), $\mathbf{D}$ is diagonal and the components $d_{i}$ are strictly positive.
+$\mathbf{W}$ is the standard Wiener process.
+The quadratic nonlinearities are energy conserving satisfying the condition $\mathbf{x}\cdot\mathbf{B}(\mathbf{x},\mathbf{x})=0$.
+This property explicitly imposes the constraint $B_{1}+B_{2}+B_{3}=0$ on the nonlinear coefficients.
+The model can be naturally cast into the general stochastic differential equation form introduced in Eq. (9):
+$\begin{bmatrix}dx_{1}\\ dx_{2}\\ dx_{3}\end{bmatrix}$ $=\Big(\underbrace{\begin{bmatrix}F_{1}\\ F_{2}\\ F_{3}\end{bmatrix}}_{\mathbf{F}}+\underbrace{\begin{bmatrix}-d_{1}&-L_{3}&L_{2}\\ L_{3}&-d_{2}&-L_{1}\\ -L_{2}&L_{1}&-d_{3}\end{bmatrix}}_{\mathbf{A}}\begin{bmatrix}x_{1}\\ x_{2}\\ x_{3}\end{bmatrix}+\underbrace{\begin{bmatrix}0&B_{1}x_{3}&0\\ -B_{1}x_{3}&0&-B_{3}x_{1}\\ 0&B_{3}x_{1}&0\end{bmatrix}}_{\mathbf{S}(\mathbf{x})}\begin{bmatrix}x_{1}\\ x_{2}\\ x_{3}\end{bmatrix}\Big)dt+\underbrace{\begin{bmatrix}\sigma_{1}&0&0\\ 0&\sigma_{2}&0\\ 0&0&\sigma_{3}\end{bmatrix}}_{\mathbf{\Sigma}}\begin{bmatrix}dW_{1}\\ dW_{2}\\ dW_{3}\end{bmatrix}$, (23)
+with $B_{2}=-(B_{1}+B_{3})$. $\mathbf{S}(\mathbf{x})$ is skew-symmetric, and therefore energy conserving.
+Performance of the approximation in a regime with dual energy cascade.
+We now test the relevance of the splitting procedure in Eq.
+(16) in the stochastic context.
+We follow Majda and Qi [58] and consider a set of parameters leading to a dual energy cascade across modes $x_{1}$, $x_{2}$, and $x_{3}$.
+Specifically, we set $d_{1}=1$, $d_{2}=d_{3}=2$; $\sigma_{1}^{2}=10$, $\sigma_{2}^{2}=\sigma_{3}^{2}=0.01$.
+The nonlinear coefficients are $B_{1}=2$, $B_{2}=B_{3}=-1$.
+The linear interactions are set to $L_{1}=0.09$, $L_{2}=0.06$, and $L_{3}=-0.03$.
+The deterministic forcing is applied only to the $x_{2}$ and $x_{3}$ modes, such that $F_{1}=0$, $F_{2}=-1$, and $F_{3}=1$.
+Given the chosen parameters, the $x_{1}$ mode is subject to strong stochastic forcing, while $x_{2}$ and $x_{3}$ are less energetic.
+The quadratic nonlinear coupling $\mathbf{B}(\mathbf{x},\mathbf{x})$ redistributes this energy, creating a cascade from the highly energetic $x_{1}$ mode to the less energetic $x_{2}$ and $x_{3}$ modes.
+Conversely, the deterministic forcing applied to $x_{2}$ and $x_{3}$ drives a backward energy cascade towards $x_{1}$.
+Therefore, these parameters place the system in a complex, dual energy cascade regime.
+We simulate the dynamics using the proposed splitting procedure for stochastic systems as in Eq.
+(16) and compare it to a standard Euler-Maruyama scheme.
+The simulation length is $5\times 10^{7}$ time steps with a step size of $\Delta t=0.001$.
+Importantly, the effective deterministic forcing $\mathbf{F}_{\text{discrete}}$ and the effective stochastic forcing $\mathbf{\Sigma}_{\text{discrete}}\bm{\xi}_{t}$ in Eq.
+(16) are derived following the steps proposed in Section II.1: $\mathbf{F}_{\text{discrete}}=\mathbf{A}^{-1}(\mathbf{M}-\mathbf{I})\mathbf{F}$, and the covariance for $\mathbf{\Sigma}_{\text{discrete}}$ is obtained by solving the associated continuous Lyapunov equation.
+In Figure 2, we show the stationary distributions of the $x_{1}$, $x_{2}$, and $x_{3}$ modes, plotted against a Gaussian distribution with the same mean and variance for reference.
+The two simulations reconstruct the identical stationary statistics and are both in good agreement with the results of Majda and Qi (see Figure 4.1C in Ref.
+[58]).
+This result further motivates Eq.
+(16) as a robust scheme to reproduce trustworthy statistics in the continuous limit $\Delta t\rightarrow 0$ beyond deterministic systems, successfully extending to forced stochastic systems.
+In the next section, we return to the data-driven setting and demonstrate that learning neural models constrained by the formulation in Eq.
+(16) yields skillful emulators even from severely sub-sampled data, a regime where both standard numerical integrators and (physics constrained) continuous-time data-driven models fail.
+Figure 2: Stationary distributions (PDFs) of $x_{1}$, $x_{2}$, and $x_{3}$ in the triad model, simulated by a standard Euler-Maruyama scheme and our Lie-Trotter splitting procedure $\mathbf{x}_{t+1}=\mathbf{F}_{\text{discrete}}+\mathbf{M}\mathbf{Q}(\mathbf{x}_{t})\mathbf{x}_{t}+\mathbf{\Sigma}_{\text{discrete}}\bm{\xi}_{t}$.
+The deterministic and stochastic forcings, $\mathbf{F}_{\text{discrete}}$ and $\mathbf{\Sigma}_{\text{discrete}}\bm{\xi}_{t}$, are derived analytically via the steps proposed in Section II.1.
+PDFs are plotted on a logarithmic scale.
+A Gaussian distribution with the same mean and variance as the Euler-Maruyama results is provided for reference.
+II.2 Data-driven emulation of the Charney-DeVore model from subsampled data
+We return to the Charney-DeVore model [13] introduced in Eq.
+(5) of the main text.
+In the main manuscript, the emulator was learned from a well-sampled trajectory to allow for a direct statistical comparison with the continuous model.
+However, our proposed neural model can operate in sparse observational regimes where standard numerical integrators fail.
+To demonstrate this, we evaluate our architecture in two coarse-grained regimes:
+  • Subsampling every 100 time steps.
+We consider a long integration ($T=10^{7}$ time steps) $\mathbf{x}_{t}\in\mathbb{R}^{n,T}$ of the Charney-DeVore model as in the main text.
+This integration was generated with an Euler-Maruyama scheme and $\Delta t=0.01$.
+We then subsample the data by considering every 100th time step.
+This corresponds to observing the data every $\Delta t=1$.
+We refer to this new subsampled trajectory as $\tilde{\mathbf{x}}_{t}$.
+We train the proposed discrete neural emulator in Eq.
+(3), with energy-conserving nonlinearities, from $\tilde{\mathbf{x}}_{t}$.
+We then assess the stationary distributions (PDFs) and autocorrelation functions (ACFs) of variables $\tilde{x}^{(1)}_{t},\tilde{x}^{(2)}_{t},\dots,\tilde{x}^{(6)}_{t}$ and show the results in Figure 3.
+Our discrete emulator remains strictly stable and successfully learns the effective discrete mapping.
+On the other hand, the numerical integrator blows up with $\Delta t=1$.
+Furthermore, data-driven models with physics constrained defined in the continuous limit also become unstable.
+  • Subsampling every 500 time steps.
+We further challenge the framework by subsampling the original time series every 500 steps.
+This reduces the amount of available data for training to 20,000 time points.
+Despite this, the PDFs and ACFs of the learned discrete mapping remain generally well-approximated (Figure 4).
+The largest error is found for the PDFs of the $x_{2}$ degree of freedom.
+Most importantly, the neural emulator guarantees stability, providing a robust finite-time flow map in a regime where traditional integration schemes fail.
+Figure 3: Experiment with subsampling size 100.
+In the first and second rows, we show the stationary distributions (PDFs) of the variables $\tilde{x}^{(1)}_{t},\tilde{x}^{(2)}_{t},\dots,\tilde{x}^{(n)}_{t}$ of the physics-constrained emulator and the subsampled time series $\tilde{\mathbf{x}}_{t}$.
+In the third and fourth rows, same as the first two rows but for the autocorrelation functions (ACFs).
+Note: a numerical solution at this coarse-grained temporal resolution does not exist (the numerical model blows up at $\Delta t=1$).
+Figure 4: Same as Fig.
+3 but after subsampling the original Charney-DeVore integration every 500 time steps (corresponding to $dt=5$).
+Note: a numerical solution at this coarse-grained temporal resolution does not exist (the numerical model blows-up at $dt=5$).
+III Fluctuation-Dissipation Theorem (FDT) from data: quasi-Gaussian and score matching
 The FDT.
-Consider a $n$-dimensional system $\mathbf{x}(t)=[x_{1},x_{2},...,x_{n}](t)$.
-The Fluctuation-Dissipation Theorem (FDT) states that the time-dependent response of an observable $A(x_{k}(t))$ to a small impulse perturbation $\delta x_{j}(0)$ imposed on $x_{j}(0)$ at time $t=0$, can be retrieved solely from stationary statistics of the system as:
-$R_{k,j}(t)=\lim_{\delta x_{j}(0)\to 0}\frac{\delta\langle A(x_{k}(t))\rangle}{\delta x_{j}(0)}=-\Big\langle x_{k}(t)\frac{\partial\ln\rho(\mathbf{x})}{\partial x_{j}}\Big|_{\mathbf{x}(0)}\Big\rangle.$ (1)
+Consider a $n$-dimensional system $\mathbf{x}_{t}=(x^{(1)}_{t},x^{(2)}_{t},...,x^{(n)}_{t})$.
+The Fluctuation-Dissipation Theorem (FDT) states that the time-dependent response of an observable $A(x^{(k)}_{t})$ to a small impulse perturbation $\delta x^{(j)}_{0}$ imposed on $x^{(j)}_{t=0}$ at time $t=0$, can be retrieved solely from stationary statistics of the system as:
+$R^{k,j}_{t}=\lim_{\delta x^{(j)}_{0}\to 0}\frac{\delta\langle A(x^{(k)}_{t})\rangle}{\delta x^{(j)}_{0}}=-\Big\langle A(x^{(k)}_{t})\frac{\partial\ln\rho(\mathbf{x})}{\partial x^{(j)}}\Big|_{\mathbf{x}_{0}}\Big\rangle.$ (24)
 where $\rho(\mathbf{x})$ represents the invariant probability distribution of the system [24, 66].
@@
 The brackets $\langle\cdot\rangle$ represents ensemble averages, in practice computed using temporal averages under the assumption of ergodicity.
-Application of the FDT in the main text focused on the observable $A(x_{k}(t))=x_{k}(t)$.
+Application of the FDT in the main text focused on the observable $A(x^{(k)}_{t})=x^{(k)}_{t}$.
 In this case, Eq.
 2 represents the impulse response in the ensemble mean.
-I.1 The FDT from data
+III.1 The FDT from data
 Quasi-Gaussian (qG) approximation.
@@
 Second, the inferred score inevitably depends on the chosen neural-network architecture and its associated hyperparameters.
-Nevertheless, the theoretical formalism of the FDT, provides analytical constraints on the instantaneous response $\mathbf{R}(t=0)$.
+Nevertheless, the theoretical formalism of the FDT, provides analytical constraints on the instantaneous response $\mathbf{R}_{t=0}$.
 As shown in [31], these constraints can be used as a posteriori correction of the estimated score, thereby ensuring reliable results.
@@
 If the interest is in the response in ensemble mean, i.e.
-observable $A(x_{k}(t))=x_{k}(t)$ in Eq.
+observable $A(x^{(k)}_{t})=x^{(k)}_{t}$ in Eq.
 (24), then it is possible to use the correction already outlined in [31].
 The response operator in the ensemble mean at time $t=0$ of any system is:
-$\mathbf{R}(0)=\mathbf{I}$ (4)
-$\mathbf{I}$ being the identity matrix. The empirical estimation of the instantaneous response operator with the inferred score $s_{\epsilon}(\mathbf{x})$ will lead to small errors, so that
-$\mathbf{R}_{\epsilon}(0)=-\frac{1}{T}\mathbf{x}^{\text{T}}\cdot s_{\epsilon}(\mathbf{x})=\mathbf{I}-\boldsymbol{\epsilon}$ (5)
-where $\mathbf{R}_{\epsilon}(0)$ denotes the estimated response. It is then possible to correct the inferred score as:
-$s_{c}(\mathbf{x})=s_{\epsilon}(\mathbf{x})\cdot(\mathbf{I}-\boldsymbol{\epsilon})^{-1}.$ (6)
+$\mathbf{R}_{0}=\mathbf{I}$ (27)
+$\mathbf{I}$ being the identity matrix. The empirical estimation of the instantaneous response operator with the inferred score $\tilde{s}(\mathbf{x})$ will lead to small errors, so that
+$\tilde{\mathbf{R}}_{0}=-\frac{1}{T}\mathbf{x}^{\text{T}}\cdot\tilde{s}(\mathbf{x})=\mathbf{I}-\boldsymbol{\epsilon}$ (28)
+where $\tilde{\mathbf{R}}_{0}$ denotes the estimated response. It is then possible to correct the inferred score as:
+$s_{c}(\mathbf{x})=\tilde{s}(\mathbf{x})\cdot(\mathbf{I}-\boldsymbol{\epsilon})^{-1}.$ (29)
 The corrected score $s_{c}(\mathbf{x})$ defined in Eq.
@@
 Note that while this correction is non-unique, different versions lead to small differences as argued in the SM of [30].
+Crucially, depending on the data, it may be useful to compute the response $R^{k,j}_{t}$ on standardized data.
+This is especially true when variables span disparate scales, as in principal component representations.
 Response operator via score matching: response in the ensemble mean.
@@
 The FDT inferred through score matching is generally more accurate than the quasi-Gaussian approximation.
-Figure 2: Time-dependent, responses in ensemble mean $x_{k}(t)$ to an impulse perturbation in the CdV model (Eq.
+Figure 5: Time-dependent, responses in ensemble mean $x^{(k)}_{t}$ to an impulse perturbation in the CdV model (Eq.
 (3) in the main text) as predicted by the FDT using the quasi-Gaussian and score-matching approximations.
-Example: column (1), row (2) quantifies the time-dependent mean response of $x_{k=2}(t)$ given a small, impulse perturbation imposed on $x_{j=1}(0)$ at time $t=0$.
-II Possible limitations of the causal constraints and practical considerations
+Example: column (1), row (2) quantifies the time-dependent mean response of $x^{k=2}_{t}$ given a small, impulse perturbation imposed on $x^{j=1}_{0}$ at time $t=0$.
+IV Possible limitations of the causal constraints and practical considerations
 Causal relations in high-dimensional stochastic dynamical systems are most naturally, and practically, identified at the level of ensemble-averaged quantities [23].
-Specifically, the response operator at the shortest time-scale can be written in terms of the ensemble averaged Jacobian as $\mathbf{R}(dt)=\langle\mathbf{J}(\mathbf{x})\rangle dt+\mathbf{I}$, where $\mathbf{J}(\mathbf{x})$ denotes the Jacobian, $\mathbf{I}$ is the identity matrix, and $dt\rightarrow 0$.
-Accordingly, the causal constraints introduced in the main text, exploit vanishing entries of the ensemble-averaged Jacobian, $\langle\partial f_{k}(\mathbf{x})/\partial x_{j}\rangle=0$, to suppress the corresponding state-dependent couplings $\partial f_{k}(\mathbf{x})/\partial x_{j}$ in the neural network loss.
+Accordingly, the causal constraints introduced in the main text, exploit vanishing entries of the ensemble-averaged Jacobian, $\langle\partial f_{k}(\mathbf{x})/\partial x^{(j)}\rangle=0$, to suppress the corresponding state-dependent couplings $\partial f_{k}(\mathbf{x})/\partial x^{(j)}$ in the neural network loss.
 A complication can in principle arise in idealized systems with exact symmetries, such as the Lorenz ’63 and Lorenz ’96 systems [50, 51].
-In such cases $R_{k,j}(dt)=0$ can arise from exact cancellations in the ensemble average, even though $\partial f_{k}(\mathbf{x})/\partial x_{j}\neq 0$ pointwise.
+In such cases $R^{k,j}_{1}=0$ can arise from exact cancellations in the ensemble average, even though $\partial f_{k}(\mathbf{x})/\partial x^{(j)}\neq 0$ pointwise.
 One possible strategy in such idealized settings is to explicitly reduce the symmetries of the underlying dynamical system, following symmetry-reduction approaches advocated in [18].
@@
 [68] proposed a nonlinear coordinate transformation to quotient the $Z_{2}$ symmetry of the Lorenz ’63 system [50], leading to the so-called “proto-Lorenz” system.
-In the case of the Lorenz ’96 system, defined as $\dot{u_{j}}=(u_{j+1}-u_{j-2})u_{j-1}-u_{j}+F,~j=0,...,J$ with periodic boundary conditions, the model is translationally invariant, with each $u_{j}$ having the same ensemble mean $\langle u_{j}\rangle=\bar{u}$.
-In this latter case, introducing a site-dependent forcing $F_{j}$ is sufficient to guarantee the validity of our framework.
-For example, one could define $F_{j}=F+\eta_{j}$, with $\eta_{j}\sim\mathcal{N}(0,\sigma^{2})$, mimicking the effect of topography, and therefore leading to a more realistic forcing term.
-Crucially, we argue that the limitation discussed above does not constitute a practical obstacle for reduced-order modeling of realistic, complex turbulent dynamical systems, with no symmetries.
-As discussed in [29], in the case of climate modeling, the effective drift $\mathbf{f}$ inferred from data is inherently non-unique and depends on modeling choices such as variable selection, resolution, filtering, and projection onto latent representations.
-In this setting, there is no single “true” dynamics; instead, models are evaluated by their ability to reproduce stationary and perturbed statistics at the chosen level of coarse-graining.
-From this perspective, a vanishing short-time response $R_{k,j}(dt)=0$ provides sufficient justification to construct an effective model
-$d\mathbf{x}=\mathbf{f}(\mathbf{x})dt+\mathbf{\Sigma}(\mathbf{x})d\mathbf{W}$
-in which the structural constraint $\partial f_{k}(\mathbf{x})/\partial x_{j}=0$ is explicitly enforced.
-The validity of such constraints is then assessed by the accuracy of the resulting stationary and forced responses, rather than by comparison to a non-existent unique underlying dynamics.
+The limitation discussed above does not constitute a practical obstacle for reduced-order modeling of realistic, complex turbulent dynamical systems.
+In these realistic flows, symmetries are always broken.
 Assessment of causal constraints.
-Finally, causally constrained emulators can be readily compared with an unconstrained counterpart by evaluating their representation of the linear response operator against the FDT-based estimate in Eq.
-(1) of the main text.
+Crucially, as also noted in the main text, differently from the rigid (data independent) physical constraints, the utility of causal constraints depends on the system under study and on the accuracy of the data-driven causal discovery.
+The benefit of augmenting physics-constrained models with causal information can be assessed directly from data by comparing the FDT-based operator with the emulators’ response operators.
 Such comparisons should be performed at short time lags to avoid spurious responses that may arise at longer lags when using the FDT [22].
@@
 In systems with exact symmetries, as discussed above, any biases introduced by the constraints will be immediately apparent in the representation of stationary variability.
-III Details on the emulator formulations
-We report the main technical details for fitting the emulators in the main text.
-This Section follows Appendix C in [29] and we refer to that work for more details.
-Consider a $n$-dimensional dynamical system described by a stochastic differential equation $d\mathbf{x}=\mathbf{f}(\mathbf{x})dt+\mathbf{\Sigma}d\mathbf{W}$.
-This equation can be discretized by the forward Euler method as follows:
-[…]
-where $\bm{\xi}(t)$ is a Gaussian white noise with zero mean, unit variance, and $\sqrt{\Delta t}\bm{\xi}(t)$ is the increment of the Wiener process $\Delta\mathbf{W}=\mathbf{W}(t+\Delta t)-\mathbf{W}(t)$.
-Without loss of generality, we rescale time units by the time sampling interval such that a non-dimensional time increment is $\Delta t=1$ [54, 53, 51, 99].
-The strategy is then to fit a rescaled version of Eq.
-(7) from observational data alone.
-Such rescaled equation takes the following form:
-$\begin{split}\Delta\mathbf{x}=\mathbf{x}(t+1)-\mathbf{x}(t)=\mathbf{f}(\mathbf{x}(t))+\mathbf{\Sigma}\bm{\xi}(t).\end{split}$ (8)
-In what follows, we will not distinguish between the original model terms (e.g., $\mathbf{f}$) and their emulator counterparts, in order to avoid unnecessary notation and maintain a clear and concise presentation.
-We explicitly clarify the distinction whenever it becomes relevant.
-We proceed in two steps: fitting the determistic drift and then the stochastic noise.
-We first focus on fitting the drift.
-We compute $\Delta\mathbf{x}=\mathbf{x}(t+1)-\mathbf{x}(t)$ and fit the deterministic dynamics such that $\Delta\mathbf{x}\approx\mathbf{f}(\mathbf{x}(t))$.
-In the main paper, we considered two cases:
-  1. i.
-    Unconstrained drift. The vanilla emulator is trained to minimize the MSE between $\Delta\mathbf{x}$ and $\mathbf{f}(\mathbf{x}(t))$ as:
-    $\mathcal{L}=MSE(\Delta\mathbf{x},\mathbf{f}(\mathbf{x}))~.$ (9)
-    The model $\mathbf{f}$ is fitted by a multilayer perceptrong (MLP).
-The MSE is minimized using a version of the gradient descent algorithm (Adam, [48]) with weight decay.
-The MLP considered here always consisted of one hidden layer with 100 neurons.
-We found empirically that adding more layers can lead to instabilities even in stationary runs, while opting for a shallow, but wide network led to stable simulations.
-The activation function used is the swish function (SiLU), which we found yields smoother responses to external perturbations.
-  2. ii.
-    Causally constrained drift. Here we consider the same architecture of the unconstrained drift above. The main difference is in the loss function, here equal to:
-    $\mathcal{L}=MSE(\Delta\mathbf{x},\mathbf{f}(\mathbf{x}))+\lambda\sum_{k,j\in\mathcal{S}}(\frac{\partial f_{k}}{\partial x_{j}})^{2}~,$ (10)
-    where the set of entries $(k,j)$ is determined by the causal Adjacency matrix introduced in the main text.
-The derivative $\partial f_{k}/\partial x_{j}$ are computed with Autograd from PyTorch, providing automatic differentiation [80].
-$\lambda$ is here set to $\lambda=10$.
-Once a determistic drift $\mathbf{f}$ is computed we close the equation with a Wiener process.
-The first step is to compute the residuals of the deterministic dynamics as $\mathbf{r}(t)=\Delta\mathbf{x}-\mathbf{f}(\mathbf{x}(t))$, $\mathbf{r}(t)\in\mathbb{R}^{n}$.
-A stochastic parametrization for this residual term is then constructed.
-As the main paper is focused on proposing constraints acting on the drift term, we adopted the simplest option, in which the noise term in Eq.
-(8) is diagonal, with each entry $(i,i)$ given by the standard deviation of $r_{i}(t)$.
-For details on more complex stochastic closure, such as state-dependent noise we refer to the work of Falasca (2025) [29].
-IV Charney De-Vore model: parameters
+V Charney De-Vore model: parameters
 The values of the parameters used for the Charney-DeVore model (Eq.
@@
 dtdt $0.01$ Integration time step
-V Practical estimation of the response operator
+VI Practical estimation of the response operator
 To estimate the response operator in practice we follow the steps proposed in Appendix A of Falasca (2025) [23]. We report those steps here for completeness.
-Given a model $d\mathbf{x}=\mathbf{f}(\mathbf{x})dt+\mathbf{\Sigma}d\mathbf{W}$, we build the impulse response operator as follows:
+Given a model $\mathbf{x}_{t+1}=\mathbf{f}(\mathbf{x}_{t})+\mathbf{\Sigma}\bm{\xi}(t)$, we build the impulse response operator as follows:
   1. 1.
@@
   2. 2.
-    For each one of the $N_{e}$ initial conditions, we impose an impulse perturbation $\Delta_{j}$ to the degree of freedom $x_{j}$ at time $t=0$.
+    For each one of the $N_{e}$ initial conditions, we impose an impulse perturbation $\Delta_{j}$ to the degree of freedom $x^{(j)}$ at time $t=0$.
 The amplitude of the perturbation should be theoretically infinitesimally small, ensuring linearity of the response even in nonlinear systems.
-In practice we do as follows: we consider the long time series of $x_{j}(t)$ from the long control integration above and define $\Delta_{j}=10^{-2}\sigma_{j}$, where $\sigma_{j}$ is the standard deviation of time series $x_{j}(t)$.
+In practice we do as follows: we consider the long time series of $x^{(j)}_{t}$ from the long control integration above and define $\Delta_{j}=10^{-2}\sigma_{j}$, where $\sigma_{j}$ is the standard deviation of time series $x^{(j)}_{t}$.
   3. 3.
-    Therefore, for a given initial condition $\mathbf{x}(0)$, we simulate two trajectories: a control trajectory without perturbation and a perturbed one, where an impulse perturbation $\Delta_{j}$ has been imposed on the $j$-th degree of freedom at time $t=0$.
+    Therefore, for a given initial condition $\mathbf{x}_{0}$, we simulate two trajectories: a control trajectory without perturbation and a perturbed one, where an impulse perturbation $\Delta_{j}$ has been imposed on the $j$-th degree of freedom at time $t=0$.
 This procedure is repeated in parallel for all initial conditions, resulting in an ensemble of $N_{e}$ paired of control and perturbed trajectories.
   4. 4.
-    At each time $t$ we then estimate the time dependent ensemble average $\langle A(x_{k}(t))\rangle$ of an observable $A(x_{k}(t))$ for both the perturbed and unperturbed run.
-We refer to the perturbed and control ensemble average respectively as $\langle A(x_{k}(t))\rangle_{\text{p}}$ and $\langle A(x_{k}(t))\rangle$.
-The observables considered in this study are going to be: (i) $A(x_{k}(t))=x_{k}(t)$ and (ii) $A(x_{k}(t))=(x_{k}(t)-\mu_{k}(t))^{2}$, with $\bm{\mu}(t)$ representing the time-dependent mean of the distribution, therefore quantifying the response in ensemble mean and variance.
+    At each time $t$ we then estimate the time dependent ensemble average $\langle A(x^{(k)}_{t})\rangle$ of an observable $A(x^{(k)}_{t})$ for both the perturbed and unperturbed run.
+We refer to the perturbed and control ensemble average respectively as $\langle A(x^{(k)}_{t})\rangle_{\text{p}}$ and $\langle A(x^{(k)}_{t})\rangle$.
+The observables considered in this study are going to be: (i) $A(x^{(k)}_{t})=x^{(k)}_{t}$ and (ii) $A(x^{(k)}_{t})=(x^{(k)}_{t}-\mu^{(k)}_{t})^{2}$, with $\bm{\mu}_{t}$ representing the time-dependent mean of the distribution, therefore quantifying the response in ensemble mean and variance.
   5. 5.
-    We define the impulse response operator for observable $A(x_{k}(t))$ as
-    $R_{k,j}(t)=\frac{\langle A(x_{k}(t))\rangle_{\text{p}}-\langle A(x_{k}(t))\rangle}{\Delta_{j}}.$
-VI Neural emulators of the Charney-DeVore model: stationary statistics
+    We define the impulse response operator for observable $A(x^{(k)}_{t})$ as
+    $R^{k,j}_{t}=\frac{\langle A(x^{(k)}_{t})\rangle_{\text{p}}-\langle A(x^{(k)}_{t})\rangle}{\Delta_{j}}.$
+VII Neural emulators of the Charney-DeVore model: stationary statistics
 Stationary density and autocorrelation functions of the Charney-deVore system as given by the numerical model, the vanilla emulator and the causal emulator.
 Figure 6: Stationary statistics.
-First and second rows: stationary distributions (PDFs) of the variables $x_{1},...,x_{6}$ of the CdV model.
-Labels: “CdV” refers to the numerical model; “Model vanilla” is the vanilla, unconstrained emulator; “Model causal” is the causally constrained emulator.
+First and second rows: stationary distributions (PDFs) of the variables $x^{(1)},...,x^{(6)}$ of the CdV model.
+Labels: “CdV” refers to the numerical model; “Physics” is the model with only physics constrained; “Physics &\& Causal” is the model constrained by both physics and causal information.
 Third and fourth rows: same as the first row but for the autocorrelation functions (ACFs).
-VII Response operator for the Charney-DeVore model
+VIII Response operator for the Charney-DeVore model
 In Figure 7 and Figure 8, we report the estimation of the full response operator in ensemble mean and variance as obtained by the numerical model and the two emulators.
@@
 This analysis is shown in Figure 9.
-Figure 4: Time-dependent, responses in ensemble mean of variable $x_{k}(t)$ to an impulse perturbation the CdV model in Eq.
+Figure 7: Time-dependent, responses in ensemble mean of variable $x^{(k)}_{t}$ to an impulse perturbation the CdV model in Eq.
 (3) in the main text as predicted by the unconstrained and causally constrained emulators.
-Example: column (1), row (2) represents what is the time-dependent response of $x_{k=2}(t)$ given a small, impulse perturbation imposed on $x_{j=1}(0)$ at time $t=0$.
-Labels: “CdV” refers to the numerical model; “Model vanilla” is the vanilla, unconstrained emulator; “Model causal” is the causally constrained emulator.
-Figure 5: Time-dependent, responses in ensemble variance of variable $x_{k}(t)$ to an impulse perturbation the CdV model in Eq.
+Example: column (1), row (2) represents what is the time-dependent response of $x^{(k=2)}_{t}$ given a small, impulse perturbation imposed on $x^{(j=1)}_{0}$ at time $t=0$.
+Labels: “CdV” refers to the numerical model; “Physics” is the model with only physics constrained; “Physics &\& Causal” is the model constrained by both physics and causal information.
+Figure 8: Time-dependent, responses in ensemble variance of variable $x^{(k)}_{t}$ to an impulse perturbation the CdV model in Eq.
 (3) in the main text as predicted by the unconstrained and causally constrained emulators.
-Example: column (1), row (2) represents what is the time-dependent response of $x_{k=2}(t)$ given a small, impulse perturbation imposed on $x_{j=1}(0)$ at time $t=0$.
-Labels: “CdV” refers to the numerical model; “Model vanilla” is the vanilla, unconstrained emulator; “Model causal” is the causally constrained emulator.
+Example: column (1), row (2) represents what is the time-dependent response of $x^{(k=2)}_{t}$ given a small, impulse perturbation imposed on $x^{(j=1)}_{0}$ at time $t=0$.
+Labels: “CdV” refers to the numerical model; “Physics” is the model with only physics constrained; “Physics &\& Causal” is the model constrained by both physics and causal information.
 Figure 9: Perturbed statistics in the linear regime.
-Panel (a): Total MSE over time for the mean response, computed from the full response operator $\mathbf{R}(t)\in\mathbb{R}^{6,6}$ for each model.
+Panel (a): Total MSE over time for the mean response, computed from the full response operator $\mathbf{R}_{t}\in\mathbb{R}^{6,6}$ for each model.
 Panel (b): Same as Panel (a) but the response in ensemble variance.
-VIII Response to large step function forcings in the Charney-DeVore model
+IX Response to large step function forcings in the Charney-DeVore model
 We now consider the nonlinear response regime by adding a step function forcing $\mathbf{F}$ to both the numerical (CdV) and neural models.
@@
 11 (variance).
-The causally coarse-grained emulator outperforms the unconstrained, vanilla emulator.
-Notably, the vanilla emulator becomes unstable in three out of six forcing experiments (rows 1, 2, and 6 in Fig.
-7), whereas the causally constrained emulator remains stable in all cases.
-Errors in the ensemble-mean response are primarily observed when the step forcing is applied to $x_{5}$ and $x_{6}$.
-To elucidate the origin of these errors, we revisit the response operator $R_{k,j}(t)$ inferred in the previous sections and focus on the ground-truth response estimated from the numerical model.
-We quantify the cumulative sensitivity of the system to perturbations applied to each degree of freedom $x_{j}$ as $D_{j}=\sum_{t=0}^{\tau_{\infty}}\sum_{k}R_{k,j}(t)$,; where is $\tau_{\infty}=40$ Model Time Units, consistent with all response analyses.
-The resulting sensitivities are shown in Fig.
-9.
-This analysis identifies $x_{5}$ and $x_{6}$ as the most sensitive degrees of freedom: perturbations applied to these variables induce the largest cumulative system-wide response.
-This high sensitivity provides a plausible, qualitative explanation for the observed degradation in ensemble-mean responses under large forcings applied to $x_{5}$ and $x_{6}$, where even small modeling errors are strongly amplified.
+The emulator constrained by both physics and causality generally leads to a better representation of the responses.
+This is clearly shown in the total MSE in Figure 12, aggregated over all response and perturbations (12), where the “physics and causal” emulator shows lower error in the long time response.
 Figure 10: Perturbed statistics in the nonlinear regime.
@@
 Row (2): same as Row (1) but using $\mathbf{F}=(0,\sigma_{2},0,0,0,0)$.
-Labels: “CdV” refers to the numerical model; “Model vanilla” is the vanilla, unconstrained emulator; “Model causal” is the causally constrained emulator.
+Labels: “CdV” refers to the numerical model; “Physics” is the model with only physics constrained; “Physics &\& Causal” is the model constrained by both physics and causal information.
 Figure 11: Perturbed statistics in the nonlinear regime.
@@
 Responses are computed using an ensemble of $N_{e}=10^{5}$ members.
-Labels: “CdV” refers to the numerical model; “Model vanilla” is the vanilla, unconstrained emulator; “Model causal” is the causally constrained emulator.
-Figure 9: Cumulative (causal) sensitivity $D_{j}=\sum_{t=0}^{\tau_{\infty}}\sum_{k}R_{k,j}(t)$ of the Charney-DeVore model in Eq.
-(3) in the main text.
-For the response operator $R_{k,j}(t)$ we use the correct ground-truth.
-The system is more sensitive to perturbations on degrees of freedom $x_{5}$ and $x_{6}$, therefore providing a possible explanation for the degradation of results when adding large, step function forcings on these degrees of freedoms.
+Labels: “CdV” refers to the numerical model; “Physics” is the model with only physics constrained; “Physics &\& Causal” is the model constrained by both physics and causal information.
+Figure 12: Perturbed statistics in the nonlinear regime. Panel (a): Total MSE over time for the ensemble mean response. Panel (b): Same as Panel (a) but the response in ensemble variance.
+X Symmetry-Broken Lorenz-96 (L96) system
+Random forcing $F_{j}$.
+In this work, we introduce a time-independent forcing in the L96 system defined as $F_{j}=F+\eta_{j}$, where $\eta_{j}\sim\mathcal{N}(0,\sigma_{F}^{2})$ with $\sigma_{F}=10$.
+Due to the random component, we report the specific realization used in our experiments for reproducibility: $F_{j}=$ (6.1087865 , 12.32213349, 28.87925261, 17.93974419, 25.202309 , 21.77103791, 9.63536354, 21.4195222 , 12.83404549, 12.77610884, 16.97167319, 4.
+, 27.92166104, 9.28910325, 26.0026942 , 17.36321124, 31.3203308 , 9.40030586, 12.88205144, 19.37769127).
+X.1 Stationary statistics
+The marginals of the invariant distribution of the L96 system predicted by the two emulators are shown in Fig. 13. The corresponding autocorrelation functions are reported in Fig. 14.
+Figure 13: Stationary statistics.
+Stationary distributions (PDFs) of the variables $x_{1},...,x_{20}$ of the L96 system comsidered in the main text.
+Labels: “L96” refers to the numerical model; “Physics” is the model with only physics constrained; “Physics &\& Causal” is the model constrained by both physics and causal information.
+Figure 14: Autocorrelation functions (ACFs) of the variables $x_{1},...,x_{20}$ of the L96 system comsidered in the main text.
+Labels: “L96” refers to the numerical model; “Physics” is the model with only physics constrained; “Physics &\& Causal” is the model constrained by both physics and causal information.